### v1

# CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering

Abstract
Retrieval-augmented generation (RAG) is widely used to ground Large Language Models (LLMs) for multi-hop question answering. Recent work mainly focused on improving answer accuracy via fine-tuning and structured or reinforcement-based optimization. However, reliable reasoning in response generation faces three challenges: 1) Reasoning Collapse. Reasoning in multi-hop QA is inherently complex due to multi-hop composition and is further destabilized by noisy retrieval. 2) Reasoning–answer inconsistency. Due to the intrinsic uncertainty of LLM generation and exposure to evidence–distractor mixtures, models may produce correct answers that are not faithfully supported by their intermediate reasoning or evidence. 3) Loss of format control. Traditional chain-of-thought generation often deviates from required structured output formats, leading to incomplete or malformed structured content. To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a Group Relative Policy Optimization (GRPO) based reinforcement learning framework that trains models to perform faithful reasoning during response generation. CRAFT employs dual reward mechanisms to optimize multi-hop reasoning: deterministic rewards ensure structural correctness while judge-based rewards verify semantic faithfulness. This optimization framework supports controllable trace variants that enable systematic analysis of how structure and scale affect reasoning performance and faithfulness. Experiments on three multi-hop QA benchmarks show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales, with the CRAFT model achieving competitive performance with closed-source LLMs across multiple reasoning trace settings.
1 Introduction
Retrieval-augmented generation (RAG) is widely used for multi-hop question answering (QA), where large language models (LLMs) generate answers grounded in retrieved candidate documents. A typical multi-hop QA task requires synthesizing information across multiple pieces of evidence that may be temporally, conceptually, or causally linked, while filtering out irrelevant or misleading distractors introduced by retrieval, calling for explicit reasoning traces that faithfully connect the supporting evidence to the final answer. As illustrated in Figure 1, reliable response generation in RAG systems faces three intertwined difficulties. First, multi-hop reasoning is inherently difficult, as it requires composing information across multiple dependent steps, and this difficulty is further amplified by noisy retrieval that introduces misleading or irrelevant documents. Second, generation by large language models is intrinsically uncertain, and under mixed evidence–distractor contexts, models may additionally exploit shortcut behaviors, producing answers that are weakly supported by the underlying reasoning or evidence. Third, models often fail to reliably follow required output formats during generation, leading to incomplete or malformed structured outputs and limiting controllability of the produced reasoning traces. These issues complicate the learning and assessment of robust reasoning in RAG systems.
Researchers have explored various approaches to improve accuracy in LLM-based multi-hop QA tasks, which can be broadly categorized along three complementary dimensions. (1) Explicit reasoning without guaranteed faithfulness. Early approaches encourage step-by-step reasoning through supervised fine-tuning (SFT), chain-of-thought prompting, and decomposition-based formulations Wei et al. (2022); Khot et al. (2023); Zhou et al. (2023); Yao et al. (2022). While effective in improving answer accuracy, the resulting reasoning traces are not guaranteed to be grounded in the supporting evidence, especially under noisy retrieval, where models may bypass relevant documents and exploit spurious shortcuts. Trivedi et al. (2020); Press et al. (2023); Lanham et al. (2023); Turpin et al. (2023). (2) Post-hoc grounding without structured reasoning. Another line of work promotes evidence grounding by introducing citations, evidence constraints, or verification signals Gao et al. (2023); Wei et al. (2025); Li et al. (2025a); Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025). But these approaches often apply grounding after generation or only at the answer level, leaving intermediate reasoning traces free-form and unstructured. (3) Optimization of LLMs with coarse-grained rewards. More recent approaches explore reinforcement learning to optimize reasoning and answer generation in retrieval-augmented QA systems Asai et al. (2024); Li et al. (2025b); Chen et al. (2025). These methods typically rely on coarse outcome-level rewards tied to final answer correctness, providing limited supervision over intermediate reasoning and faithfulness properties, and allowing correct answers to be produced without aligned or faithful reasoning.
To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) for the response generation stage of RAG-based LLMs in multi-hop QA under noisy retrieved candidates. CRAFT trains generators to produce structured, machine-auditable reasoning traces, enabling systematic verification of reasoning faithfulness to retrieved evidence and generated answers. Our study consists of two aspects: 1) RL training with fine-grained reward supervision. Given a question and a retrieved candidate set containing both evidence and distractors, CRAFT optimizes reasoning trace generation using group-based reinforcement learning. Training is guided by a decomposed reward function that combines (i) deterministic rewards enforcing machine-checkable constraints, such as output structure validity, citation-set correctness when applicable, and answer matching, with (ii) judge rewards from a high-capacity LLM-as-a-judge that assess semantic faithfulness, including consistency across reasoning stages (e.g., planreasonanswer) and grounding to retrieved evidence. This design provides targeted supervision over unlabeled but critical semantic properties of multi-hop reasoning. 2) Controllable trace space and scale study. CRAFT defines a controllable space of reasoning trace variants by composing trace components (<plan>, <gold_docs>, <reason>, <answer>) along an auditability spectrum. This design enables systematic analysis of how trace structure influences performance and faithfulness, as well as how model scale interacts with trace complexity and audit constraints. Our contributions are three-fold:
Dual reward mechanisms for faithful multi-hop reasoning optimization. We propose CRAFT, a GRPO-based RL framework combining deterministic rewards for verifiable metrics and judge rewards for semantic faithfulness.
Systematic analysis of controllable reasoning traces. We study how different combinations of trace components affect model performance, and characterize how model scale shapes the preferred trace structures and their performance boundaries in small LLMs.
Experimental results demonstrate that CRAFT consistently improves answer accuracy and the faithfulness of generated reasoning traces across model scales. Notably, the CRAFT-trained 7B model achieves competitive performance with closed-source LLMs across different reasoning trace configurations on three multi-hop QA benchmarks.
2 Related Works
2.1 Faithful Reasoning in LLMs
Chain-of-thought (CoT) prompting Wei et al. (2022) improves reasoning in language models, but recent work questions whether these traces faithfully reflect the model’s actual decision process. Lanham et al. (2023) found that CoT faithfulness decreases with model scale, while Turpin et al. (2023) showed that CoT can be influenced by spurious input features; in safety-critical LLM-integrated systems, such unconstrained outputs undermine response reliability Zhang et al. (2025). Recent evaluation methods include Tutek et al. (2025)’s faithfulness measurement via unlearning reasoning steps, Arakelyan et al. (2025)’s FLARE framework using logic programming for verifiable reasoning, and Sui et al. (2025)’s approach anchoring responses to structured knowledge graph paths. However, these methods largely rely on post hoc analysis rather than enforcing machine-checkable structure during training.
2.2 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) Lewis et al. (2020) grounds language models in retrieved documents. For multi-hop reasoning, Trivedi et al. (2023) proposed IRCoT with interleaved retrieval, while recent work includes HopRAG Liu et al. (2025a) using passage graphs, DualRAG Cheng et al. (2025) combining reasoning-augmented querying with knowledge aggregation, and OPERA Liu et al. (2025b) coordinating planning and retrieval via RL. On generation, Gao et al. (2023) trained models for citations, while InstructRAG Wei et al. (2025) and RAG-DDR Li et al. (2025a) optimize generation with reward signals. Despite these advances, most approaches prioritize retrieval quality or answer-level grounding rather than producing structured reasoning traces.
2.3 Reinforcement Learning for LLMs
Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) aligns language models via policy optimization like PPO Schulman et al. (2017), DPO Rafailov et al. (2023), and RRHF Yuan et al. (2023). GRPO Shao et al. (2024) removes the need for a critic by normalizing rewards within sampled groups, enabling efficient training for reasoning models like DeepSeek-R1 Guo et al. (2025). For RAG, Chen et al. (2025) applied multi-agent RL, while R3-RAG Li et al. (2025b) jointly optimizes reasoning and retrieval with outcome and process rewards. However, most methods use scalar rewards, making it hard to provide targeted supervision over structured reasoning.
2.4 Multi-hop Question Answering
Multi-hop QA benchmarks include HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). Press et al. (2023) identified a compositionality gap in models, while Trivedi et al. (2020) showed they exploit shortcuts rather than genuine reasoning. Methods like Decomposed Prompting Khot et al. (2023), Least-to-Most Zhou et al. (2023), and ReAct Yao et al. (2022) improve performance. Recent work has focused on understanding reasoning capabilities: Yao et al. (2025) found that models can learn implicit multi-hop reasoning but require exponentially more training data as reasoning depth increases, while Yu et al. (2025) proposed back attention to enhance latent multi-hop reasoning. For evaluation, Lee et al. (2025) introduced GRADE, which generates difficulty-controlled QA pairs based on reasoning depth and retrieval difficulty. Despite these advances, most approaches lack machine-checkable grounding to verify that reasoning steps are supported by evidence.
3 Method
3.1 Problem Formulation
We study the response generation stage of RAG-based LLMs for multi-hop QA tasks, where each instance consists of a question and a retrieved candidate document set that mixes supporting evidence with distractors. Answering typically requires reasoning over multiple supporting documents indexed by , while distractors introduce spurious shortcuts. Our goal is to train models that generate not only correct answers but also auditable reasoning traces—structured outputs in XML-style format (<plan>, <gold_docs>, <reason>, <answer>) that expose the derivation process. Formally, given and a gold answer , the model outputs a trace containing components (plan), (citation set), (reasoning), and (answer), satisfying: (i) format parsability , (ii) answer correctness , and (iii) trace faithfulness , where enforces internal consistency and evidence grounding. This formulation transforms the learning problem from maximizing a sparse answer reward to optimizing a composite objective over the structure, process, and outcome of reasoning.
3.2 Overview
As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) a CRAFT Interface that defines structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
CRAFT Interface.
We define a family of XML-based output templates spanning an auditability spectrum. As detailed in Algorithm 1 (upper part), the full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning. Intermediate variants (CRAFT–CRAFT) selectively ablate <plan> or <gold_docs>, while CRAFT (Answer-Only) serves as a minimal baseline. This design enforces a chain of faithfulness: —any hallucination breaks at least one link and is detectable.
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team (2024) | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
| Qwen2.5-1.5B Team (2024) | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
| Qwen2.5-3B Team (2024) | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
| Qwen2.5-7B Team (2024) | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
| Qwen3-4B Team (2025) | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
| Qwen3-30B Team (2025) | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
| API | ||||||||||
| GPT-5-mini OpenAI (2025) | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
| Deepseek-V3.2 DeepSeek-AI (2025) | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
| Gemini-2.5-Flash Google DeepMind (2025) | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team (2024) | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
| Qwen2.5-1.5B Team (2024) | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
| Qwen2.5-3B Team (2024) | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
| Qwen2.5-7B Team (2024) | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 7.25 | 13.13 | 0.80 | 24.80 | 34.61 | 5.40 | 22.85 | 28.46 | 1.85 |
| CRAFT | Trained | 19.35 | 27.99 | 28.75 | 45.70 | 60.01 | 57.25 | 44.70 | 52.05 | 47.85 |
| CRAFT | Trained | 30.96 | 39.69 | 45.85 | 54.53 | 69.06 | 79.50 | 56.50 | 64.75 | 76.10 |
| CRAFT | Trained | 54.35 | 65.03 | 80.35 | 66.61 | 82.09 | 95.00 | 78.72 | 84.57 | 94.80 |
3.3 CRAFT Training
We align models using GRPO with dual reward mechanisms to maximize a composite objective that calibrates structure, process, and outcome.
Principle
We optimize the policy to generate traces that are not only correct but structurally valid and reason-faithful, effectively supervising the “chain of faithfulness”. For each input , GRPO samples a group of traces and normalizes rewards within the group:
where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing . The key design choice is the total reward , which we decompose into four components:
where weights control the trade-off between strictness and accuracy.
3.3.1 (1) Deterministic Rewards
Rewards of this type can be computed directly from the trace structure and gold labels without requiring an LLM judge.
Format Compliance ().
We enforce strict adherence to the XML schema to ensure parsability.
where is the set of strings containing all required tags for variant in correct order. We set all other reward terms to when .
Citation Reward ().
We encourage document-level citations that cover the supporting documents while avoiding distractors. When <gold_docs> is present, we parse the predicted citation set and compute an index-level F1 against the gold supports :
Answer Correctness ().
We target final task accuracy with soft-match supervision.
where parses the <answer> content from the trace and is the standard token-level F1 score after normalization. This term ensures that process constraints do not degrade final performance, providing a continuous gradient toward the gold answer.
3.3.2 (2) Judge Rewards
Rewards of this type rely on an LLM judge to assess semantic consistency and evidence grounding, supervising the reasoning process rather than only the outcome.
Faithfulness Audit ().
We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1 for the detailed procedure). Let denote extracting the trace fields from . We employ a high-capacity LLM judge to compute:
by averaging multiple machine-checkable criteria:
where and depends on the trace variant (e.g., plan/citation checks apply only when the corresponding fields exist). Concretely, the judge checks: (i) planreason (), (ii) citation-setreason (), (iii) reasonanswer (), and (iv) grounding (), i.e., key claims in are supported by the cited document text in .
Why no direct supervision for Plan/Reason text?
Unlike evidence selection (discrete indices) or final answers (short spans), the intermediate <plan> and <reason> are open-ended natural language generations with high semantic variance. Direct supervision (e.g., via SFT on teacher traces) often leads to brittle imitation or format collapse on out-of-distribution data. Instead of enforcing a rigid “gold plan,” our rewards focus on functional correctness via consistency checks (): a plan is valid if it is followed; a reason is valid if it is grounded. This allows the model to discover its own optimal reasoning paths within the auditable structure.
4 Experiment
| Model | Version | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Qwen2.5-7B | ||||||||||
| CRAFT | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 | |
| CRAFT | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 | |
| CRAFT | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 | |
| CRAFT | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 | |
| CRAFT | 31.55 | 41.97 | - | 53.40 | 67.92 | - | 50.60 | 59.29 | - | |
| Qwen2.5-7B | ||||||||||
| CRAFT | 32.35(-4.5) | 43.38(-4.4) | 50.85(-7.4) | 57.75(+0.8) | 72.04(+0.5) | 83.50(-0.2) | 62.65(+4.95) | 71.32(+4.6) | 78.95(+2.1) | |
| CRAFT | 29.85(-4.5) | 41.17(-3.5) | 56.10(-9.4) | 55.70(+0.9) | 70.25(+0.5) | 86.70(-1.4) | 59.25(+3.2) | 67.73(+2.1) | 78.30(+1.1) | |
| CRAFT | 35.60(-3.4) | 47.66(-1.9) | 61.35(-7.4) | 59.25(+1.5) | 73.68(+1.6) | 88.25(-2.4) | 65.40(+4.75) | 74.33(+4.5) | 85.80(+1.6) | |
| CRAFT | 31.80(-6.6) | 42.71(-6.6) | 66.60(-10.0) | 57.90(-1.4) | 72.43(-2.0) | 89.35(-4.4) | 63.15(+2.4) | 72.00(+1.8) | 82.25(-3.9) | |
| CRAFT | 27.00(-4.6) | 37.82(-4.2) | - | 53.30(-0.1) | 68.12(+0.2) | - | 56.15(+5.6) | 63.79(+4.5) | - | |
| CRAFT w/o judge reward | ||||||||||
| CRAFT | 51.37(+14.5) | 62.09(+14.3) | 71.20(+13.0) | 63.62(+6.7) | 79.40(+7.9) | 91.80(+8.1) | 75.91(+18.2) | 81.96(+15.2) | 89.25(+12.4) | |
| CRAFT | 55.93(+21.6) | 66.00(+21.4) | 69.45(+4.0) | 64.70(+9.9) | 80.01(+10.3) | 93.20(+5.2) | 77.45(+21.4) | 83.47(+17.8) | 84.80(+7.6) | |
| CRAFT | 52.75(+13.8) | 62.71(+13.2) | 72.15(+3.5) | 64.18(+6.4) | 79.21(+7.1) | 93.80(+3.1) | 74.98(+14.3) | 81.48(+11.7) | 88.00(+3.8) | |
| CRAFT | 55.82(+17.5) | 65.24(+15.9) | 76.85(+0.2) | 65.15(+5.9) | 80.97(+6.6) | 95.20(+1.5) | 78.09(+17.3) | 83.76(+13.6) | 88.80(+2.6) | |
| CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
| CRAFT (full) | ||||||||||
| CRAFT | 54.35(+17.5) | 65.03(+17.2) | 80.35(+22.1) | 66.61(+9.7) | 82.09(+10.5) | 95.00(+11.4) | 78.72(+21.0) | 84.57(+17.9) | 94.80(+18.0) | |
| CRAFT | 58.35(+24.0) | 68.29(+23.7) | 78.35(+12.9) | 67.28(+12.5) | 82.10(+12.4) | 95.20(+7.2) | 79.70(+23.7) | 85.56(+19.9) | 89.90(+12.7) | |
| CRAFT | 54.65(+15.7) | 64.59(+15.1) | 81.40(+12.7) | 66.09(+8.3) | 81.04(+8.9) | 95.70(+5.1) | 76.88(+16.2) | 83.34(+13.5) | 93.20(+9.0) | |
| CRAFT | 56.91(+18.6) | 66.47(+17.2) | 86.70(+10.1) | 66.51(+7.2) | 81.91(+7.5) | 96.00(+2.3) | 79.39(+18.6) | 85.14(+14.9) | 94.10(+8.0) | |
| CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
4.1 Experimental Setup
Datasets.
We evaluate CRAFT on three multi-hop QA benchmarks: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions Press et al. (2023); Trivedi et al. (2020). For training, we construct a mixed dataset by sampling from all three benchmarks, comprising 20,000 entries in total. For evaluation, we test on 2,000 samples from each benchmark.
Baselines.
We compare against two model categories: (1) Open-source pre-trained models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) evaluated via in-context prompting; (2) Closed-source API models: GPT OpenAI (2025), DeepSeek DeepSeek-AI (2025), and Gemini Google DeepMind (2025). For our method, we train models at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO Shao et al. (2024).
Metrics.
We report three evaluation metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes the token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (as detailed in Algorithm 1), providing a complementary measure beyond answer correctness.
Implementation Details.
All experiments are conducted on 4 NVIDIA H100 GPUs. We use the Qwen2.5 series as base models and train with GRPO using decomposed rewards. Training hyperparameters and configurations are detailed in Appendix B.
4.2 Main Results
Table 1 presents results evaluating CRAFT against open-source pre-trained models, closed-source API models, and SFT baselines across three multi-hop QA benchmarks. Additional experiments in Appendix C.
CRAFT improves accuracy across models at all scales.
For the 7B model, CRAFT achieves gains of +17.5%/+9.7%/+21.0% EM on MuSiQue/HotpotQA/2WikiMHQA compared to base Qwen2.5-7B, corresponding to relative improvements of 47%/17%/36%. The transformation is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model—a 16 improvement, showing how CRAFT’s decomposed rewards unlock latent reasoning capacity in small models.
CRAFT improves faithfulness across models at all scales.
For the 7B model, CRAFT achieves +22.0%/+8.9%/+16.1% EM improvements over SFT on MuSiQue/HotpotQA/2WikiMHQA. While SFT often degrades Faithfulness compared to base models (50.85% vs. 58.25% for 7B on MuSiQue, a drop of -7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT. This divergence illustrates that SFT learns to imitate surface patterns, whereas CRAFT’s judge-based faithfulness reward encourages internally consistent, evidence-grounded traces.
CRAFT matches or exceeds closed-source API models.
54.35% EM on MuSiQue (vs. GPT-5-mini 53.45%), 66.61%/82.09% EM/F1 on HotpotQA (vs. 61.30%/80.09%), and 78.72% EM on 2WikiMHQA (vs. 70.55%, +8.2%). This demonstrates that GRPO can train a 7B open-source model to match or exceed proprietary systems.
4.3 Ablation Study
We conduct two ablation studies: (1) trace template ablation with aligned deterministic rewards, and (2) judge reward ablation. In CRAFT, the trace template determines deterministic reward availability (, , ), so variants ablate both trace components and their corresponding rewards. Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>, disabling all deterministic rewards and judge reward. We further ablate the judge reward by training with and without . Table 2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
Removing structure degrades performance.
For base models, answer-only CRAFT achieves the lowest EM (31.55%/53.40%/50.60% on MuSiQue/HotpotQA/2WikiMHQA), underperforming CRAFT (38.35%/59.30%/60.80%) by 5.9%–10.2%. While GRPO improves CRAFT substantially to 56.25%/63.25%/67.15%, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%). Notably, CRAFT cannot be evaluated for Faithfulness as it lacks reasoning traces.
Trace components reveal accuracy-faithfulness trade-offs.
For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without explicit training. After GRPO training, this gap narrows substantially (CRAFT: 80.35%/95.00%/94.80% vs. CRAFT: 86.70%/96.00%/94.10%), while CRAFTv1/CRAFTv2 improve auditability via explicit planning and citation.
CRAFT outperforms SFT across all configurations.
CRAFT achieves gains across all variants: +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models. In contrast, SFT shows frequent degradation (up to -6.6% EM on CRAFT) and systematic Faithfulness drops (-7.4% to -10.0% on MuSiQue). For CRAFT, GRPO improves MuSiQue EM by +17.5% (36.85%54.35%) while SFT decreases it by -4.5% (36.85%32.35%).
Judge reward is essential for faithfulness.
Without judge reward, Faithfulness drops -8.9% to -9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable. This validates our dual reward design: deterministic rewards ensure structural correctness and answer accuracy, while judge rewards are essential for semantic faithfulness.
4.4 Training and Traces Analysis
| Question: Which member of the research project conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005 was Kevin R. Henke most critical of? |
| Gold Answer: Russell Humphreys |
| plan |
| Subquestion1: Which research project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005? |
| Subquestion2: Who was Kevin R. Henke most critical of in [ANSWER_0]? |
| gold_docs |
| [2, 7] |
| reason |
| Subquestion1: From document [2], it is clear that the RATE project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005. |
| Subquestion2: Document [7] states that Kevin R. Henke has been critical of the RATE project’s results, specifically mentioning that he has argued that Russell Humphreys, a young-earth creationist involved in the project, has made errors in his research. |
| answer |
| Russell Humphreys ✓ |
4.4.1 Training Dynamics
Training performance scales with model size. Figure 3 shows total reward rising monotonically with scale: 0.5B (), 1.5B (), 3B (), and 7B (). Across all scales, reward curves converge smoothly without oscillation, indicating stable GRPO optimization under our multi-reward design. Notably, the largest relative gain appears at 1.5B, while marginal improvements diminish at 7B, suggesting saturation once most reward components approach their upper bounds.
Sub-rewards reveal fast format learning and a faithfulness threshold. Format Reward () converges within 50 training steps across all scales (0.5B: ; 7B remains near throughout), confirming that structural constraints are quickly internalized once the output template is learned. Answer Reward () scales with model capacity, reaching final scores of 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B, while Gold_doc Reward () improves steadily (0.050.27; 0.120.66; 0.190.55; 0.760.92), reflecting stronger evidence grounding in larger models.
Faithfulness exhibits a clear capacity threshold. Faithfulness Reward () remains near zero for 0.5B () and reaches only modest levels at 1.5B (0.10), whereas 3B and 7B show sharp improvements (, ). This transition indicates that structured consistency auditing is substantially more demanding than surface format compliance or answer correctness, and only emerges once the model has sufficient capacity to internalize and satisfy judge-based semantic constraints.
4.4.2 Output Traces Case Study
We analyze the quality of generated reasoning traces to understand how CRAFT structures multi-hop reasoning. Table 3 presents a representative example using the CRAFT, demonstrating CRAFT’s systematic approach with explicit planning, gold document citation, step-by-step reasoning, and final answer extraction. Additional case study in Appendix A. About traces format analysis in Appendix C.3.
5 Conclusion
We present CRAFT, a GRPO-based reinforcement learning framework that trains RAG-based LLMs to generate structured, machine-auditable reasoning traces for multi-hop QA. CRAFT employs a dual reward design that combines deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness. Experiments on three benchmarks show that CRAFT consistently improves both accuracy and faithfulness over base models and SFT, with CRAFT matching or exceeding closed-source API models. Ablation results highlight the importance of structured traces and judge rewards for learning faithful reasoning, while training dynamics analysis indicates that faithfulness learning requires sufficient model capacity.
Appendix A Faithfulness Case Study
We present two representative cases from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs, demonstrating how CRAFT training improves both answer correctness and reasoning faithfulness. Each case includes the complete judge evaluation across all four faithfulness dimensions defined in Algorithm 1.
A.1 Case 1: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
This case demonstrates how an unfaithful and incorrect response is transformed into a faithful and correct one.
Question: Who won the Indy car race in the largest city of the state where the Oh Yeah album performer was born?
Gold Answer: Mario Andretti
| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
| Plan |
Subq1: Where is the Oh Yeah album from?
Subq2: What is the largest city and capital of that state? Subq3: Who won the Indy car race in that city? |
Subq1: Where is the Oh Yeah performer from?
Subq2: What is the largest city and capital of the state where the Oh Yeah performer is from? Subq3: Who won the Indy car race in the largest city and capital of that state? |
| Gold Docs | [5, 12, 14] | [11, 12, 1] |
| Reason | [Answer1] Charles Mingus, the creator of the Oh Yeah album, was born in Nogales, Arizona (doc [5]). [Answer2] The largest city and capital of Arizona is Phoenix (doc [12]). [Answer3] The Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and the winner of the Indy car race there in 2016 was Ryan Hunter-Reay (doc [1]). | Doc [11] indicates that the Oh Yeah performer, Charles Mingus, is from Nogales, Arizona. Doc [12] states that the capital and largest city of Arizona is Phoenix. Doc [1] mentions that the Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and it is an Indy car race where Mario Andretti won. |
| Answer | Ryan Hunter-Reay | Mario Andretti ✓ |
| Judge Evaluation | ||
| Plan-Reason () | 1 - Plan steps followed in order | 1 - Plan followed: performer→state→city→winner |
| Receipt-Reason () | 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs [11,12,1] in gold_docs |
| Reason-Answer () | 1 - Answer matches reasoning conclusion | 1 - “Mario Andretti” supported by reasoning |
| Evidence Grounded () | 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified in cited docs |
| Overall () | 0.0 - Failures: doc_outside_gold_docs, evidence_not_supported | 1.0 - All checks passed |
A.2 Case 2: Base (EM=1, Faith=0) CRAFT (EM=1, Faith=1)
This case demonstrates that correct answers can arise from unfaithful reasoning, and CRAFT enforces faithfulness even when accuracy is preserved.
Question: What is the average yearly income for a person of the same nationality as the creator of MacGruder and Loud?
Gold Answer: $72,641
| Aspect | Base Model (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
| Plan |
Subq1: Who is the creator of MacGruder and Loud?
Subq2: What is the nationality of [Answer1]? Subq3: What is the average yearly income for a person of [Answer2] nationality? |
Subq1: Who is the creator of MacGruder and Loud?
Subq2: What is the average yearly income for a person of the same nationality as the creator? |
| Gold Docs | [4, 5, 2] | [8, 4] |
| Reason | [Answer1]: From document 4, we learn that the creator of MacGruder and Loud is Aaron Spelling. Document 5 tells us that Lorraine Crosby, known as “Mrs. Loud,” is English. Therefore, the nationality of the creator of MacGruder and Loud is English. Document 2 provides the average yearly income for Americans, which is the closest nationality we have data for. The mean household income in the United States was $72,641. | Doc [8] states that MacGruder and Loud was one of the few failures from Aaron Spelling’s production company. Doc [4] provides information about Aaron Spelling, indicating that he was an American film and television producer. Therefore, the creator of MacGruder and Loud is American. The average yearly income for an American, as mentioned in doc [2], is $72,641. |
| Answer | $72,641 ✓ (correct answer, but unfaithful reasoning) | $72,641 ✓ |
| Judge Evaluation | ||
| Plan-Reason () | 0 - Identifies Aaron Spelling but concludes “English” | 1 - Plan followed: creator→nationality→income |
| Receipt-Reason () | 0 - Misuses doc [5] (Lorraine Crosby creator) | 1 - All cited docs [8,4] in gold_docs |
| Reason-Answer () | 0 - Derives “English” but uses US income | 1 - “$72,641” supported by reasoning |
| Evidence Grounded () | 0 - “Creator is English” unsupported | 1 - Spelling→American→US income verified |
| Overall () | 0.0 - Failures: plan_ignored, answer_not_supported, evidence_not_supported | 1.0 - All checks passed |
A.3 Summary
These cases illustrate two critical aspects of CRAFT’s faithfulness enforcement:
Case 1 shows that CRAFT corrects both unfaithful reasoning (citing documents outside gold_docs, hallucinating unsupported claims) and incorrect answers.
Case 2 demonstrates that answer correctness does not imply reasoning faithfulness. The Base model produces the correct answer through a flawed reasoning chain (confusing Lorraine Crosby with Aaron Spelling), while CRAFT ensures the answer is derived through a consistent, evidence-grounded process.
The judge evaluation tables reveal that CRAFT training optimizes all four faithfulness dimensions simultaneously: plan-reason consistency (), receipt-reason consistency (), reason-answer consistency (), and evidence grounding (). This multi-dimensional supervision prevents shortcut learning and ensures that models produce not just correct answers, but trustworthy reasoning traces.
Appendix B Implementation Details
Training Framework.
We implement CRAFT using MS-Swift Zhao et al. [2025], an efficient and scalable fine-tuning framework for large language models. MS-Swift provides native support for GRPO training with vLLM acceleration, enabling efficient on-policy sampling during reinforcement learning. All experiments are conducted on NVIDIA H100 96GB x4 GPUs.
Model Configuration.
We train four model scales based on the Qwen2.5-Instruct series: 0.5B, 1.5B, 3B, and 7B parameters. All models use bfloat16 precision with Flash Attention 2 for memory-efficient training. Table 6 summarizes the key hyperparameters.
| Hyperparameter | Value |
| GRPO Configuration | |
| Batch Generations Accum. | 4 4 64 |
| Effective Batch Size | 1,024 |
| Max Completion Length | 256 tokens |
| Optimization | |
| Learning Rate | 5e-6 |
| LR Scheduler | Cosine |
| Warmup Ratio | 0.1 |
| Infrastructure | |
| Precision | bfloat16 |
| Attention | Flash Attention 2 |
| vLLM Max Length | 16,384 tokens |
Judge Model.
For the faithfulness reward and evaluation, we employ Qwen3-Max Team [2025] as the LLM judge. This model provides reliable semantic consistency assessment while maintaining computational efficiency. Because Qwen3-Max already functions as the judge for reward computation and evaluation, it is excluded from all performance comparison experiments to avoid role overlap.
Reward Function Configuration.
The reward functions vary across prompt templates based on their structural components:
CRAFT/CRAFT: , , (includes <gold_docs> field)
CRAFT/CRAFT: , (no explicit document citation)
CRAFT: , (direct answer generation)
All reward weights are set to 1.0, treating each component equally. The faithfulness reward is computed via LLM judge during evaluation but not used as a training signal in the ablation experiments (Table 2, “w/o judge reward”).
Training Data.
Each prompt template uses a dedicated training dataset of 20,000 samples, formatted according to the template’s structural requirements. The data distribution is: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA samples. Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns that indicate diverse linguistic characteristics across the three multi-hop QA datasets.
Appendix C Additional Experiments
This section presents comprehensive results for prompt template variants CRAFT–CRAFT, including training dynamics and evaluation results across all model scales.
C.1 Training Dynamics
Figures 5–8 illustrate the training dynamics for CRAFT–CRAFT templates. Consistent with CRAFT (main paper Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
Template-specific patterns: (1) CRAFT (Figure 5) shows similar dynamics to CRAFT, with all four reward components (, , , ) contributing to optimization. (2) CRAFT/CRAFT (Figures 6–7) lack but maintain stable training through format and answer rewards. (3) CRAFT (Figure 8) uses only and —notably, the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
Faithfulness reward verification.
To verify that the low training faithfulness reward (Figure 3) reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics (Algorithm 1) on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0). Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%. This indicates that while 69.7% of correct predictions exhibit fully faithful reasoning, 30.3% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness. This test-time value is approximately six times higher than the training-time average of 0.11 observed for this model (Figure 3), confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, resulting in lower average rewards during training, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
| Qwen3-4B Team [2025] | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
| Qwen3-30B Team [2025] | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
| Qwen2.5-1.5B Team [2024] | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
| Qwen2.5-3B Team [2024] | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
| Qwen2.5-7B Team [2024] | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 6.50 | 12.76 | 9.85 | 20.60 | 29.86 | 23.70 | 22.65 | 28.79 | 15.90 |
| CRAFT | Trained | 19.26 | 27.66 | 26.67 | 45.27 | 59.20 | 51.64 | 45.45 | 52.40 | 38.78 |
| CRAFT | Trained | 32.05 | 41.48 | 59.40 | 55.69 | 70.06 | 89.80 | 54.70 | 63.17 | 78.90 |
| CRAFT | Trained | 58.35 | 68.29 | 78.35 | 67.28 | 82.10 | 95.20 | 79.70 | 85.56 | 89.90 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
| Qwen3-4B Team [2025] | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
| Qwen3-30B Team [2025] | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
| Qwen2.5-1.5B Team [2024] | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
| Qwen2.5-3B Team [2024] | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
| Qwen2.5-7B Team [2024] | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 4.75 | 10.09 | 7.95 | 20.20 | 30.14 | 20.85 | 20.50 | 25.65 | 12.50 |
| CRAFT | Trained | 18.30 | 26.89 | 22.57 | 44.18 | 57.84 | 55.25 | 42.86 | 48.34 | 42.29 |
| CRAFT | Trained | 34.11 | 43.71 | 61.20 | 57.24 | 71.15 | 92.60 | 54.90 | 63.47 | 80.10 |
| CRAFT | Trained | 54.65 | 64.59 | 81.40 | 66.09 | 81.04 | 95.70 | 76.88 | 83.34 | 93.20 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
| Qwen3-4B Team [2025] | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
| Qwen3-30B Team [2025] | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
| Qwen2.5-1.5B Team [2024] | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
| Qwen2.5-3B Team [2024] | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
| Qwen2.5-7B Team [2024] | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
| Ours (RL) | ||||||||||
| CRAFT | Trained | 8.45 | 15.11 | 25.40 | 25.95 | 36.20 | 40.70 | 24.70 | 30.94 | 26.50 |
| CRAFT | Trained | 16.41 | 23.27 | 29.57 | 43.68 | 57.79 | 65.66 | 44.22 | 51.33 | 48.96 |
| CRAFT | Trained | 32.71 | 43.14 | 71.40 | 56.11 | 70.63 | 94.30 | 55.68 | 64.18 | 82.00 |
| CRAFT | Trained | 56.91 | 66.47 | 86.70 | 66.51 | 81.91 | 96.00 | 79.39 | 85.14 | 94.10 |
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | |||
| EM | F1 | EM | F1 | EM | F1 | ||
| Open-source | |||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
| Qwen3-4B Team [2025] | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
| Qwen3-30B Team [2025] | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
| API | |||||||
| GPT-5-mini OpenAI [2025] | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
| SFT | |||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
| Qwen2.5-1.5B Team [2024] | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
| Qwen2.5-3B Team [2024] | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
| Qwen2.5-7B Team [2024] | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
| Ours (RL) | |||||||
| CRAFT | Trained | 0.00 | 0.00 | 0.00 | 0.00 | 0.05 | 0.05 |
| CRAFT | Trained | 22.95 | 31.91 | 45.70 | 59.04 | 38.40 | 46.19 |
| CRAFT | Trained | 31.80 | 42.07 | 53.15 | 67.10 | 49.30 | 57.19 |
| CRAFT | Trained | 56.25 | 65.38 | 63.25 | 78.00 | 67.15 | 73.77 |
| Model | Type | CRAFT | CRAFT | CRAFT | CRAFT | CRAFT | |||||
| Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | ||
| Qwen2.5-0.5B | |||||||||||
| Avg. | 0.0 | 98.3(+98.3) | 1.1 | 99.2(+98.1) | 0.0 | 94.3(+94.3) | 1.0 | 99.4(+98.4) | 9.0 | 0.1(-8.9) | |
| HotpotQA | 0.0 | 99.6(+99.6) | 0.1 | 99.8(+99.7) | 0.1 | 97.5(+97.4) | 0.8 | 99.8(+99.0) | 5.7 | 0.1(-5.6) | |
| 2WikiMQA | 0.0 | 99.5(+99.5) | 3.2 | 99.7(+96.5) | 0.0 | 92.5(+92.5) | 2.3 | 99.8(+97.5) | 16.5 | 0.1(-16.4) | |
| MuSiQue | 0.0 | 95.8(+95.8) | 0.0 | 98.2(+98.2) | 0.0 | 92.8(+92.8) | 0.0 | 98.5(+98.5) | 4.9 | 0.0(-4.9) | |
| Qwen2.5-1.5B | |||||||||||
| Avg. | 0.0 | 99.3(+99.3) | 2.1 | 95.1(+93.0) | 0.2 | 99.5(+99.3) | 1.2 | 81.9(+80.7) | 0.8 | 100.0(+99.2) | |
| HotpotQA | 0.0 | 99.5(+99.5) | 2.9 | 98.9(+96.0) | 0.1 | 99.2(+99.1) | 1.0 | 92.9(+91.9) | 0.7 | 100.0(+99.3) | |
| 2WikiMQA | 0.0 | 99.4(+99.4) | 0.8 | 99.2(+98.4) | 0.0 | 99.5(+99.5) | 1.4 | 91.2(+89.8) | 1.4 | 100.0(+98.6) | |
| MuSiQue | 0.1 | 99.1(+99.0) | 2.5 | 87.4(+84.9) | 0.6 | 99.7(+99.1) | 1.2 | 61.6(+60.4) | 0.2 | 99.9(+99.7) | |
| Qwen2.5-3B | |||||||||||
| Avg. | 13.2 | 96.9(+83.7) | 50.3 | 98.8(+48.5) | 6.4 | 98.0(+91.6) | 15.6 | 98.1(+82.5) | 41.6 | 100.0(+58.4) | |
| HotpotQA | 14.8 | 98.2(+83.4) | 55.2 | 99.2(+44.0) | 4.3 | 99.4(+95.1) | 27.6 | 99.2(+71.6) | 32.8 | 100.0(+67.2) | |
| 2WikiMQA | 7.7 | 99.0(+91.3) | 33.8 | 99.4(+65.6) | 6.3 | 99.1(+92.8) | 8.2 | 98.8(+90.6) | 23.9 | 100.0(+76.1) | |
| MuSiQue | 17.2 | 93.5(+76.3) | 61.8 | 97.6(+35.8) | 8.7 | 95.6(+86.9) | 10.8 | 96.3(+85.5) | 68.1 | 100.0(+31.9) | |
| Qwen2.5-7B | |||||||||||
| Avg. | 94.1 | 99.3(+5.2) | 99.2 | 100.0(+0.8) | 94.4 | 97.2(+2.8) | 61.6 | 99.8(+38.2) | 15.4 | 100.0(+84.6) | |
| HotpotQA | 96.1 | 100.0(+3.9) | 99.5 | 100.0(+0.5) | 95.7 | 99.1(+3.4) | 49.5 | 100.0(+50.5) | 11.2 | 100.0(+88.8) | |
| 2WikiMQA | 95.4 | 99.1(+3.7) | 99.3 | 100.0(+0.7) | 96.0 | 97.0(+1.0) | 72.9 | 100.0(+27.1) | 9.6 | 100.0(+90.4) | |
| MuSiQue | 90.9 | 98.9(+8.0) | 98.9 | 100.0(+1.1) | 91.6 | 95.5(+3.9) | 62.5 | 99.4(+36.9) | 25.4 | 100.0(+74.6) | |
C.2 Evaluation Results
CRAFT Results.
CRAFT retains the same four-field structure as CRAFT but with simplified instructions. CRAFT achieves strong performance: 58.35% EM on MuSiQue (+24.0% vs Base), 67.28% on HotpotQA (+12.5% vs Base), and 79.70% on 2WikiMHQA (+23.7% vs Base). Notably, CRAFT surpasses all API models on MuSiQue and 2WikiMHQA while achieving competitive Faithfulness scores (78.35%–95.20%).
CRAFT Results.
CRAFT removes explicit document citation (<gold_docs>), testing whether models can maintain faithfulness without explicit grounding signals. CRAFT achieves 54.65% EM on MuSiQue, 66.09% on HotpotQA, and 76.88% on 2WikiMHQA. Despite lacking supervision, the model maintains high Faithfulness (81.40%–95.70%), suggesting that plan-reason-answer consistency provides sufficient structure for faithful reasoning.
CRAFT Results.
CRAFT further simplifies the template while maintaining structured reasoning. CRAFT achieves the highest Faithfulness scores among all templates: 86.70% on MuSiQue, 96.00% on HotpotQA, and 94.10% on 2WikiMHQA. EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance when combined with appropriate reward design.
CRAFT Results.
CRAFT represents the minimal template with direct answer generation, lacking structured reasoning traces ( and not applicable). CRAFT achieves 56.25% EM on MuSiQue, 63.25% on HotpotQA, and 67.15% on 2WikiMHQA. A critical observation: CRAFT completely fails on this template (near-zero performance), indicating that smaller models require structured reasoning traces for stable training. This validates the importance of trace structure for model capacity constraints.
CRAFT Template Sensitivity Analysis.
Our experiments reveal a striking template sensitivity pattern unique to the smallest model, with different failure modes for SFT and GRPO training. SFT training collapses on templates CRAFT, CRAFT, and CRAFT (EM 1.5% across benchmarks), achieving meaningful performance only on CRAFT (1.10/12.80/22.45%) and CRAFT (0.50/11.75/23.15%). In contrast, GRPO rescues templates CRAFT–CRAFT but still fails on CRAFT. Specifically, GRPO achieves 7.25/24.80/22.85% on CRAFT, 6.50/20.60/22.65% on CRAFT, 4.75/20.20/20.50% on CRAFT, and 8.45/25.95/24.70% on CRAFT, while CRAFT remains completely broken (0.00/0.00/0.05%).
This dichotomy reveals two key insights: (1) SFT cold-start is fragile for 0.5B models—templates CRAFT/CRAFT with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse. (2) GRPO’s exploration mechanism overcomes SFT’s brittleness—reinforcement learning enables the model to discover valid reasoning paths even on complex structures (CRAFT/CRAFT), as long as some structural scaffolding exists. (3) Template CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure—direct answer generation lacks sufficient guidance for stable learning.
The optimal template for 0.5B deployment is CRAFT (<plan>+<reason>+<answer> without <gold_docs>), which provides the best balance: enough structure to guide reasoning without exceeding representational capacity, and compatibility with both SFT initialization and GRPO refinement.
C.3 Output Traces Format Analysis
Table 11 presents a comprehensive analysis of format compliance across all CRAFT variants and model scales.
GRPO dramatically improves format compliance across all scales. Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates (CRAFT–CRAFT), while even 3B models only reach 6.4%–50.3% on average. In stark contrast, GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases. Even 7B base models, despite already achieving 90.9%–99.5% on complex templates (CRAFT–CRAFT), benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%). This demonstrates that reward successfully teaches models to follow structured output requirements across all capacity levels.
Format complexity exhibits scale-dependent learnability. For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a puzzling anomaly on CRAFT, where GRPO models (0.1%) underperform base (9.0%) by -4.9% to -16.4% across datasets. This suggests the model defaults to more complex formats learned during multi-template training. In contrast, larger models master all formats: 1.5B achieves +80.7% to +99.3% gains, 3B reaches +58.4% to +91.6% improvements, and 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
Format Error Analysis. Table 12 presents representative error patterns across all CRAFT variants. Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
| Version | Base Model Errors (7B) | GRPO Model Errors (7B) |
| CRAFT |
Missing closing tags (60%):
<answer>1958 Award missing </answer> Truncation (30%): ...grandmother of Henry St John ( response cut mid-sentence |
Empty answer (rare):
<answer> no content, missing </answer> |
| CRAFT |
Missing closing tags (70%):
<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): <reason>...Michael Rubbo </answer> wrong closing tag None bare text |
No errors found ✓ |
| CRAFT |
Missing <answer> block (45%):
<reason>...answer is American. </reason> entire <answer>...</answer> missing Truncation (40%): ...cannot determine based on response cut off |
Missing/malformed tags (3%):
<answer>Infanta Maria missing </answer> <answer>The Woman in Fifth </ truncated closing tag |
| CRAFT |
Missing closing tags (85%):
<answer>Russell Humphreys missing </answer> |
Extra text after tags (0.2%):
<answer>No</answer> (Note: The reasoning...) parenthetical note not allowed |
| CRAFT |
Wrong format (85%):
A Bug’s Life no XML tags [Beaumont, Texas] bracket format [No] should be <answer>No</answer> |
No errors found ✓ |
C.4 Human Annotation Validation
| Aspect | Both | Both | LLM | LLM | Agr. | |
| Pass | Fail | Human | Human | |||
| Plan-Reason | 388 | 83 | 18 | 11 | 94.2% | 0.81 |
| Evidence Ground. | 380 | 80 | 23 | 17 | 92.0% | 0.75 |
| Answer Deriv. | 400 | 80 | 10 | 10 | 96.0% | 0.86 |
| Gold Doc Citation | 370 | 80 | 30 | 20 | 90.0% | 0.70 |
| Overall | 385 | 80 | 20 | 15 | 93.0% | 0.78 |
To validate the reliability of our LLM-based faithfulness metric, we conducted a human evaluation study following established protocols Adlakha et al. [2024]; Zheng et al. [2023]. Two expert annotators independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT with CRAFT template, using the same four-aspect criteria as our LLM judge (Algorithm 1). We measure inter-annotator reliability using Cohen’s Kappa Artstein [2017]:
where is the observed agreement and is the expected agreement by chance. Here denotes both-pass cases, denotes both-fail cases, (LLM pass, human fail) and (LLM fail, human pass) are disagreements, and is the total sample size. Table 13 presents the detailed per-aspect results. The results demonstrate consistently high agreement across all four aspects, with the overall faithfulness achieving 93.0% agreement and between automatic and human evaluation, indicating substantial agreement Zheng et al. [2023]; Bavaresco et al. [2024]. Answer Derivation shows the highest agreement (), as the correctness of final answers is relatively unambiguous. Gold Document Citation has lower agreement (), reflecting occasional ambiguity in determining whether cited documents are strictly necessary. Disagreements (20–50 cases per aspect) primarily occurred in edge cases involving implicit reasoning or borderline evidence usage. This strong alignment confirms that our LLM-based faithfulness metric serves as a reliable and scalable proxy for human judgment.

### v2

@@
-CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering
+CRAFT: Calibrated Reasoning with Answer-Faithful Traces for Multi-Hop Question Answering
 Abstract
-Retrieval-augmented generation (RAG) is widely used to ground Large Language Models (LLMs) for multi-hop question answering.
-Recent work mainly focused on improving answer accuracy via fine-tuning and structured or reinforcement-based optimization.
-However, reliable reasoning in response generation faces three challenges: 1) Reasoning Collapse.
-Reasoning in multi-hop QA is inherently complex due to multi-hop composition and is further destabilized by noisy retrieval.
-2) Reasoning–answer inconsistency.
-Due to the intrinsic uncertainty of LLM generation and exposure to evidence–distractor mixtures, models may produce correct answers that are not faithfully supported by their intermediate reasoning or evidence.
-3) Loss of format control.
-Traditional chain-of-thought generation often deviates from required structured output formats, leading to incomplete or malformed structured content.
-To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a Group Relative Policy Optimization (GRPO) based reinforcement learning framework that trains models to perform faithful reasoning during response generation.
-CRAFT employs dual reward mechanisms to optimize multi-hop reasoning: deterministic rewards ensure structural correctness while judge-based rewards verify semantic faithfulness.
-This optimization framework supports controllable trace variants that enable systematic analysis of how structure and scale affect reasoning performance and faithfulness.
-Experiments on three multi-hop QA benchmarks show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales, with the CRAFT model achieving competitive performance with closed-source LLMs across multiple reasoning trace settings.
+Retrieval-augmented large language models, when optimized with outcome-level rewards, can achieve strong answer accuracy on multi-hop questions.
+However, under noisy retrieval, models frequently suffer from right-answer-wrong-reason failures: they may exploit spurious shortcuts or produce reasoning traces weakly grounded in the supporting evidence.
+Furthermore, the lack of structured output control prevents reliable auditing of the underlying reasoning quality.
+To address this, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework for the response generation stage of retrieval-augmented multi-hop question answering.
+CRAFT trains models to produce structured reasoning traces with configurable levels of auditability (e.g., by selectively retaining planning, evidence citation, or reasoning steps).
+Training combines two complementary forms of supervision: deterministic rewards enforce verifiable constraints, including format compliance, answer correctness, and citation-set validity, while a judge-based reward audits semantic faithfulness by evaluating reasoning consistency and evidence grounding.
+Experiments show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales.
+Notably, semantic judge-based rewards improve answer accuracy rather than compromise it, enabling CRAFT to achieve performance competitive with strong closed-source models.
+CRAFT: Calibrated Reasoning with Answer-Faithful Traces for Multi-Hop Question Answering
+Yu Liu1, 2††thanks: Equal contribution., Wenxiao Zhang311footnotemark: 1, Diandian Guo1, 2, Cong Cao1††thanks: Corresponding authors., Fangfang Yuan1, Qiang Sun3, Yanbing Liu1, 2, Jin B.
+Hong3, Zhiyuan Ma422footnotemark: 2 1Institute of Information Engineering, CAS, 2University of Chinese Academy of Sciences, 3The University of Western Australia, 4Huazhong University of Science and Technology caocong@iie.ac.cn, mzyth@hust.edu.cn
 1 Introduction
-Retrieval-augmented generation (RAG) is widely used for multi-hop question answering (QA), where large language models (LLMs) generate answers grounded in retrieved candidate documents.
-A typical multi-hop QA task requires synthesizing information across multiple pieces of evidence that may be temporally, conceptually, or causally linked, while filtering out irrelevant or misleading distractors introduced by retrieval, calling for explicit reasoning traces that faithfully connect the supporting evidence to the final answer.
-As illustrated in Figure 1, reliable response generation in RAG systems faces three intertwined difficulties.
-First, multi-hop reasoning is inherently difficult, as it requires composing information across multiple dependent steps, and this difficulty is further amplified by noisy retrieval that introduces misleading or irrelevant documents.
-Second, generation by large language models is intrinsically uncertain, and under mixed evidence–distractor contexts, models may additionally exploit shortcut behaviors, producing answers that are weakly supported by the underlying reasoning or evidence.
-Third, models often fail to reliably follow required output formats during generation, leading to incomplete or malformed structured outputs and limiting controllability of the produced reasoning traces.
-These issues complicate the learning and assessment of robust reasoning in RAG systems.
-Researchers have explored various approaches to improve accuracy in LLM-based multi-hop QA tasks, which can be broadly categorized along three complementary dimensions.
+Retrieval-augmented generation (RAG) grounds large language models (LLMs) in external documents and is widely used for multi-hop question answering (QA), where answers require composing information across multiple passages Lewis et al.
+(2020).
+However, a correct answer does not imply faithful reasoning: models can exploit spurious shortcuts—citing irrelevant documents, confusing entities, or following logically inconsistent chains—yet still reach the gold answer (Figure 1, top).
+Such right-answer-wrong-reason failures are invisible to standard metrics that evaluate only the final output.
+The problem is especially acute in multi-hop QA, which demands compositional inference over dependent steps (e.g., resolving a bridge entity and linking its attributes to downstream facts) and is further destabilized by distractors that share surface overlap with the query Trivedi et al.
+(2020); Press et al.
+(2023).
+Prior work addresses aspects of this problem from three directions, but each leaves significant gaps.
 (1) Explicit reasoning without guaranteed faithfulness.
-Early approaches encourage step-by-step reasoning through supervised fine-tuning (SFT), chain-of-thought prompting, and decomposition-based formulations Wei et al.
+Chain-of-thought and decomposition-based methods encourage step-by-step reasoning Wei et al.
 (2022); Khot et al.
@@
 (2023); Yao et al.
-(2022).
-While effective in improving answer accuracy, the resulting reasoning traces are not guaranteed to be grounded in the supporting evidence, especially under noisy retrieval, where models may bypass relevant documents and exploit spurious shortcuts.
-Trivedi et al.
-(2020); Press et al.
-(2023); Lanham et al.
+(2022), but the resulting traces are not grounded in the retrieved evidence; under noisy retrieval, models may bypass relevant documents and produce plausible-looking but unfaithful rationales Lanham et al.
 (2023); Turpin et al.
@@
 (2) Post-hoc grounding without structured reasoning.
-Another line of work promotes evidence grounding by introducing citations, evidence constraints, or verification signals Gao et al.
+Citation and verification mechanisms introduce evidence signals such as document-level citations or consistency checks Gao et al.
 (2023); Wei et al.
@@
 (2025); Sui et al.
-(2025).
-But these approaches often apply grounding after generation or only at the answer level, leaving intermediate reasoning traces free-form and unstructured.
-(3) Optimization of LLMs with coarse-grained rewards.
-More recent approaches explore reinforcement learning to optimize reasoning and answer generation in retrieval-augmented QA systems Asai et al.
+(2025), yet these are typically applied after generation or enforced only at the answer level, leaving intermediate reasoning steps unstructured and difficult to audit at scale.
+(3) Outcome-level optimization without process supervision.
+Reinforcement learning approaches optimize reasoning and generation Asai et al.
 (2024); Li et al.
 (2025b); Chen et al.
+(2025), but most rely on coarse outcome-level rewards tied solely to final answer correctness, providing weak credit assignment for multi-hop traces and no supervision over intermediate reasoning faithfulness.
+To address these limitations, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) Shao et al.
+(2024) for the response generation stage of retrieval-augmented multi-hop QA.
+CRAFT defines a family of structured output templates that decompose generation into auditable stages (planning, evidence selection, step-by-step reasoning, and answer extraction), forming trace variants with configurable levels of auditability (Figure 1, bottom).
+Within each variant, these stages form a chain of faithfulness where each component constrains the next, so that unfaithful reasoning is more likely to violate at least one verifiable link.
+Training combines two complementary forms of supervision: deterministic rewards enforce format validity, citation-set correctness, and answer matching, while a judge-based reward from a high-capacity LLM audits semantic faithfulness by verifying cross-stage consistency (planreason, citationreason, reasonanswer) and grounding individual claims against the retrieved evidence.
+This extends process-level supervision beyond rule-based objectives to open-ended semantic properties (e.g., reasoning consistency and claim verifiability).
+The judge operates on binary decomposed criteria to limit bias propagation, and its reliability is validated against human annotations (§4.5).
+Together, these designs enable the model to learn not just what to answer but how to reason faithfully.
+Our contributions are as follows:
+We propose CRAFT, an RL framework for retrieval-augmented QA that generates controllable trace variants (CRAFT–CRAFT), improving performance and reasoning auditability.
+We introduce a dual-reward design that combines deterministic rewards for verifiable constraints with judge-based rewards for reasoning faithfulness, extending process-level supervision from rule-based objectives to open-ended semantic properties.
+Experiments across multiple model scales show that CRAFT substantially improves reasoning faithfulness after training, and further reveal that semantic judge-based rewards can improve answer accuracy rather than compromise it.
+CRAFT achieves faithfulness and accuracy comparable to larger closed-source models.
+2 Related Work
+2.1 Faithful Reasoning and Grounding in RAG
+CoT prompting Wei et al.
+(2022) improves reasoning but may not faithfully reflect the model’s decision process Lanham et al.
+(2023); Turpin et al.
+(2023); Zhang et al.
 (2025).
-These methods typically rely on coarse outcome-level rewards tied to final answer correctness, providing limited supervision over intermediate reasoning and faithfulness properties, and allowing correct answers to be produced without aligned or faithful reasoning.
-To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) for the response generation stage of RAG-based LLMs in multi-hop QA under noisy retrieved candidates.
-CRAFT trains generators to produce structured, machine-auditable reasoning traces, enabling systematic verification of reasoning faithfulness to retrieved evidence and generated answers.
-Our study consists of two aspects: 1) RL training with fine-grained reward supervision.
-Given a question and a retrieved candidate set containing both evidence and distractors, CRAFT optimizes reasoning trace generation using group-based reinforcement learning.
-Training is guided by a decomposed reward function that combines (i) deterministic rewards enforcing machine-checkable constraints, such as output structure validity, citation-set correctness when applicable, and answer matching, with (ii) judge rewards from a high-capacity LLM-as-a-judge that assess semantic faithfulness, including consistency across reasoning stages (e.g., planreasonanswer) and grounding to retrieved evidence.
-This design provides targeted supervision over unlabeled but critical semantic properties of multi-hop reasoning.
-2) Controllable trace space and scale study.
-CRAFT defines a controllable space of reasoning trace variants by composing trace components (<plan>, <gold_docs>, <reason>, <answer>) along an auditability spectrum.
-This design enables systematic analysis of how trace structure influences performance and faithfulness, as well as how model scale interacts with trace complexity and audit constraints.
-Our contributions are three-fold:
-Dual reward mechanisms for faithful multi-hop reasoning optimization.
-We propose CRAFT, a GRPO-based RL framework combining deterministic rewards for verifiable metrics and judge rewards for semantic faithfulness.
-Systematic analysis of controllable reasoning traces.
-We study how different combinations of trace components affect model performance, and characterize how model scale shapes the preferred trace structures and their performance boundaries in small LLMs.
-Experimental results demonstrate that CRAFT consistently improves answer accuracy and the faithfulness of generated reasoning traces across model scales.
-Notably, the CRAFT-trained 7B model achieves competitive performance with closed-source LLMs across different reasoning trace configurations on three multi-hop QA benchmarks.
-2 Related Works
-2.1 Faithful Reasoning in LLMs
-Chain-of-thought (CoT) prompting Wei et al.
-(2022) improves reasoning in language models, but recent work questions whether these traces faithfully reflect the model’s actual decision process.
-Lanham et al.
-(2023) found that CoT faithfulness decreases with model scale, while Turpin et al.
-(2023) showed that CoT can be influenced by spurious input features; in safety-critical LLM-integrated systems, such unconstrained outputs undermine response reliability Zhang et al.
-(2025).
-Recent evaluation methods include Tutek et al.
-(2025)’s faithfulness measurement via unlearning reasoning steps, Arakelyan et al.
-(2025)’s FLARE framework using logic programming for verifiable reasoning, and Sui et al.
-(2025)’s approach anchoring responses to structured knowledge graph paths.
-However, these methods largely rely on post hoc analysis rather than enforcing machine-checkable structure during training.
-2.2 Retrieval-Augmented Generation
-Retrieval-augmented generation (RAG) Lewis et al.
-(2020) grounds language models in retrieved documents.
-For multi-hop reasoning, Trivedi et al.
-(2023) proposed IRCoT with interleaved retrieval, while recent work includes HopRAG Liu et al.
-(2025a) using passage graphs, DualRAG Cheng et al.
-(2025) combining reasoning-augmented querying with knowledge aggregation, and OPERA Liu et al.
-(2025b) coordinating planning and retrieval via RL.
-On generation, Gao et al.
-(2023) trained models for citations, while InstructRAG Wei et al.
-(2025) and RAG-DDR Li et al.
-(2025a) optimize generation with reward signals.
-Despite these advances, most approaches prioritize retrieval quality or answer-level grounding rather than producing structured reasoning traces.
-2.3 Reinforcement Learning for LLMs
-Reinforcement learning from human feedback (RLHF) Ouyang et al.
-(2022) aligns language models via policy optimization like PPO Schulman et al.
+Post-hoc faithfulness methods Tutek et al.
+(2025); Arakelyan et al.
+(2025); Sui et al.
+(2025) do not enforce structure during training.
+RAG Lewis et al.
+(2020) grounds models in documents, with multi-hop extensions including IRCoT Trivedi et al.
+(2023), HopRAG Liu et al.
+(2025a), and OPERA Liu et al.
+(2025b).
+Generation grounding via citations Gao et al.
+(2023) or reward signals Wei et al.
+(2025); Li et al.
+(2025a) primarily targets retrieval quality or answer-level grounding rather than structured reasoning with training-time faithfulness enforcement.
+2.2 Reinforcement Learning for Reasoning
+RLHF Ouyang et al.
+(2022) aligns models via PPO Schulman et al.
 (2017), DPO Rafailov et al.
 (2023), and RRHF Yuan et al.
+(2023); GRPO Shao et al.
+(2024) removes the critic via group-wise normalization for reasoning models Guo et al.
 (2025).
-GRPO Shao et al.
-(2024) removes the need for a critic by normalizing rewards within sampled groups, enabling efficient training for reasoning models like DeepSeek-R1 Guo et al.
+Process rewards improve exploration efficiency Setlur et al.
+(2025) but require domain-specific verifier data.
+For RAG, multi-agent RL Chen et al.
+(2025) and joint reasoning-retrieval optimization Li et al.
+(2025b) have been explored, yet most methods use scalar answer-correctness rewards.
+LLM-as-a-Judge Zheng et al.
+(2023) offers scalable reward signals but introduces biases Chen et al.
+(2024); Gu et al.
+(2024); mitigations include binary criteria, rubric-based scoring, and human validation Adlakha et al.
+(2024); Bavaresco et al.
 (2024).
-For RAG, Chen et al.
-(2025) applied multi-agent RL, while R3-RAG Li et al.
-(2025b) jointly optimizes reasoning and retrieval with outcome and process rewards.
-However, most methods use scalar rewards, making it hard to provide targeted supervision over structured reasoning.
+CRAFT adopts these strategies with binary pass/fail judgments across decomposed dimensions, complemented by deterministic rewards.
 2.3 Multi-hop Question Answering
-Multi-hop QA benchmarks include HotpotQA Yang et al.
-(2018), 2WikiMultiHopQA Ho et al.
-(2020), and MuSiQue Trivedi et al.
+Multi-hop QA benchmarks Yang et al.
+(2018); Ho et al.
+(2020); Trivedi et al.
+(2022) require composing information across multiple documents, yet models exhibit a compositionality gap Press et al.
+(2023) and exploit shortcuts Trivedi et al.
 (2020).
-Press et al.
-(2023) identified a compositionality gap in models, while Trivedi et al.
-(2020) showed they exploit shortcuts rather than genuine reasoning.
-Methods like Decomposed Prompting Khot et al.
-(2023), Least-to-Most Zhou et al.
-(2023), and ReAct Yao et al.
-(2022) improve performance.
-Recent work has focused on understanding reasoning capabilities: Yao et al.
-(2025) found that models can learn implicit multi-hop reasoning but require exponentially more training data as reasoning depth increases, while Yu et al.
-(2025) proposed back attention to enhance latent multi-hop reasoning.
-For evaluation, Lee et al.
-(2025) introduced GRADE, which generates difficulty-controlled QA pairs based on reasoning depth and retrieval difficulty.
-Despite these advances, most approaches lack machine-checkable grounding to verify that reasoning steps are supported by evidence.
+Decomposition-based Khot et al.
+(2023); Zhou et al.
+(2023) and agentic Yao et al.
+(2022) methods improve performance, while recent work examines implicit reasoning Yao et al.
+(2025), back attention Yu et al.
+(2025), and difficulty-controlled evaluation Lee et al.
+(2025).
+Most approaches lack machine-checkable grounding to verify reasoning faithfulness.
 3 Method
@@
 Formally, given and a gold answer , the model outputs a trace containing components (plan), (citation set), (reasoning), and (answer), satisfying: (i) format parsability , (ii) answer correctness , and (iii) trace faithfulness , where enforces internal consistency and evidence grounding.
-This formulation transforms the learning problem from maximizing a sparse answer reward to optimizing a composite objective over the structure, process, and outcome of reasoning.
 3.2 Overview
-As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) a CRAFT Interface that defines structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
-CRAFT Interface.
+As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) CRAFT Trace Variants that define structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
+Trace Templates.
 We define a family of XML-based output templates spanning an auditability spectrum.
-As detailed in Algorithm 1 (upper part), the full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning.
-Intermediate variants (CRAFT–CRAFT) selectively ablate <plan> or <gold_docs>, while CRAFT (Answer-Only) serves as a minimal baseline.
+The full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning.
+Intermediate variants selectively ablate components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains only <reason>+<answer>, and CRAFT (Answer-Only) serves as a minimal baseline.
 This design enforces a chain of faithfulness: — any hallucination breaks at least one link and is detectable.
-| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
-| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
-| Open-source | ||||||||||
-| Qwen2.5-0.5B Team (2024) | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
-| Qwen2.5-1.5B Team (2024) | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
-| Qwen2.5-3B Team (2024) | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
-| Qwen2.5-7B Team (2024) | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
-| Qwen3-4B Team (2025) | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
-| Qwen3-30B Team (2025) | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
-| API | ||||||||||
-| GPT-5-mini OpenAI (2025) | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
-| Deepseek-V3.2 DeepSeek-AI (2025) | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
-| Gemini-2.5-Flash Google DeepMind (2025) | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
-| SFT | ||||||||||
-| Qwen2.5-0.5B Team (2024) | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
-| Qwen2.5-1.5B Team (2024) | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
-| Qwen2.5-3B Team (2024) | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
-| Qwen2.5-7B Team (2024) | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
-| Ours (RL) | ||||||||||
-| CRAFT | Trained | 7.25 | 13.13 | 0.80 | 24.80 | 34.61 | 5.40 | 22.85 | 28.46 | 1.85 |
-| CRAFT | Trained | 19.35 | 27.99 | 28.75 | 45.70 | 60.01 | 57.25 | 44.70 | 52.05 | 47.85 |
-| CRAFT | Trained | 30.96 | 39.69 | 45.85 | 54.53 | 69.06 | 79.50 | 56.50 | 64.75 | 76.10 |
-| CRAFT | Trained | 54.35 | 65.03 | 80.35 | 66.61 | 82.09 | 95.00 | 78.72 | 84.57 | 94.80 |
+The trace template also determines which rewards are applicable: variants containing <gold_docs> activate , while variants with reasoning fields activate .
 3.3 CRAFT Training
 We align models using GRPO with dual reward mechanisms to maximize a composite objective that calibrates structure, process, and outcome.
-Principle
+Principle.
 We optimize the policy to generate traces that are not only correct but structurally valid and reason-faithful, effectively supervising the “chain of faithfulness”.
 For each input , GRPO samples a group of traces and normalizes rewards within the group:
-where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing .
+where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing:
 The key design choice is the total reward , which we decompose into four components:
 where weights control the trade-off between strictness and accuracy.
-3.3.1 (1) Deterministic Rewards
+3.3.1 Deterministic Rewards
 Rewards of this type can be computed directly from the trace structure and gold labels without requiring an LLM judge.
@@
 This term ensures that process constraints do not degrade final performance, providing a continuous gradient toward the gold answer.
-3.3.2 (2) Judge Rewards
+3.3.2 Judge Reward
 Rewards of this type rely on an LLM judge to assess semantic consistency and evidence grounding, supervising the reasoning process rather than only the outcome.
 Faithfulness Audit ().
-We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1 for the detailed procedure).
-Let denote extracting the trace fields from .
-We employ a high-capacity LLM judge to compute:
+We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1). Let denote extracting the trace fields from . We employ a high-capacity LLM judge to compute:
 by averaging multiple machine-checkable criteria:
@@
 Concretely, the judge checks: (i) planreason (), (ii) citation-setreason (), (iii) reasonanswer (), and (iv) grounding (), i.e., key claims in are supported by the cited document text in .
-Why no direct supervision for Plan/Reason text?
-Unlike evidence selection (discrete indices) or final answers (short spans), the intermediate <plan> and <reason> are open-ended natural language generations with high semantic variance.
-Direct supervision (e.g., via SFT on teacher traces) often leads to brittle imitation or format collapse on out-of-distribution data.
-Instead of enforcing a rigid “gold plan,” our rewards focus on functional correctness via consistency checks (): a plan is valid if it is followed; a reason is valid if it is grounded.
-This allows the model to discover its own optimal reasoning paths within the auditable structure.
-4 Experiment
+Each check yields a binary score; is their average.
+The binary, dimension-decomposed design limits the scope for judge error compared to holistic scoring.
+The judge model (Qwen3-Max) is distinct from the trained policy to avoid self-enhancement bias.
+Its reliability is validated against human annotations (§4.5), and cross-judge evaluation with an independent model confirms that results are not specific to the choice of judge (Appendix C).
+| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
+| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
+[…]
+| Qwen2.5-0.5B | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
+| Qwen2.5-1.5B | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
+| Qwen2.5-3B | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
+| Qwen2.5-7B | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
+| Qwen3-4B | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
+| Qwen3-30B | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
+[…]
+| GPT-5-mini | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
+| Deepseek-V3.2 | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
+| Gemini-2.5-Flash | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
+[…]
+| Qwen2.5-0.5B | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
+| Qwen2.5-1.5B | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
+| Qwen2.5-3B | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
+| Qwen2.5-7B | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
+[…]
+| CRAFT | Trained | 7.25 | 13.13 | 0.80 | 24.80 | 34.61 | 5.40 | 22.85 | 28.46 | 1.85 |
+| CRAFT | Trained | 19.35 | 27.99 | 28.75 | 45.70 | 60.01 | 57.25 | 44.70 | 52.05 | 47.85 |
+| CRAFT | Trained | 30.96 | 39.69 | 45.85 | 54.53 | 69.06 | 79.50 | 56.50 | 64.75 | 76.10 |
+| CRAFT | Trained | 54.35 | 65.03 | 80.35 | 66.61 | 82.09 | 95.00 | 78.72 | 84.57 | 94.80 |
 | Model | Version | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
-| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
-| Qwen2.5-7B | ||||||||||
+| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
+[…]
 | CRAFT | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 | |
@@
 | CRAFT | 31.55 | 41.97 | - | 53.40 | 67.92 | - | 50.60 | 59.29 | - | |
-| Qwen2.5-7B | ||||||||||
+[…]
 | CRAFT | 32.35(-4.5) | 43.38(-4.4) | 50.85(-7.4) | 57.75(+0.8) | 72.04(+0.5) | 83.50(-0.2) | 62.65(+4.95) | 71.32(+4.6) | 78.95(+2.1) | |
@@
 | CRAFT | 27.00(-4.6) | 37.82(-4.2) | - | 53.30(-0.1) | 68.12(+0.2) | - | 56.15(+5.6) | 63.79(+4.5) | - | |
-| CRAFT w/o judge reward | ||||||||||
+[…]
 | CRAFT | 51.37(+14.5) | 62.09(+14.3) | 71.20(+13.0) | 63.62(+6.7) | 79.40(+7.9) | 91.80(+8.1) | 75.91(+18.2) | 81.96(+15.2) | 89.25(+12.4) | |
@@
 | CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
-| CRAFT (full) | ||||||||||
+[…]
 | CRAFT | 54.35(+17.5) | 65.03(+17.2) | 80.35(+22.1) | 66.61(+9.7) | 82.09(+10.5) | 95.00(+11.4) | 78.72(+21.0) | 84.57(+17.9) | 94.80(+18.0) | |
@@
 | CRAFT | 56.25(+24.7) | 65.38(+23.4) | - | 63.25(+9.9) | 78.00(+10.1) | - | 67.15(+16.6) | 73.77(+14.5) | - | |
+Why no direct supervision for Plan/Reason?
+Unlike (discrete indices) and (short spans), and are open-ended generations with high semantic variance, making direct supervision via teacher traces brittle and prone to format collapse.
+Instead, we supervise functional correctness: is valid iff follows it (); is valid iff it is grounded ().
+This lets the model discover optimal reasoning paths within the auditable structure without imitating a fixed trace.
+4 Experiments
 4.1 Experimental Setup
@@
 (2022).
-These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions Press et al.
-(2023); Trivedi et al.
-(2020).
-For training, we construct a mixed dataset by sampling from all three benchmarks, comprising 20,000 entries in total.
+These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions.
+For training, we construct a mixed dataset of 20,000 entries sampled from all three benchmarks.
 For evaluation, we test on 2,000 samples from each benchmark.
 Baselines.
-We compare against two model categories: (1) Open-source pre-trained models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) evaluated via in-context prompting; (2) Closed-source API models: GPT OpenAI (2025), DeepSeek DeepSeek-AI (2025), and Gemini Google DeepMind (2025).
-For our method, we train models at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO Shao et al.
-(2024).
+We compare against: (1) Open-source models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) via in-context prompting; (2) Closed-source API models: GPT-5-mini OpenAI (2025), DeepSeek-V3.2 DeepSeek-AI (2025), and Gemini-2.5-Flash Google DeepMind (2025).
+We train CRAFT at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO.
 Metrics.
-We report three evaluation metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes the token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (as detailed in Algorithm 1), providing a complementary measure beyond answer correctness.
+We report three metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (Algorithm 1).
 Implementation Details.
-All experiments are conducted on 4 NVIDIA H100 GPUs.
-We use the Qwen2.5 series as base models and train with GRPO using decomposed rewards.
-Training hyperparameters and configurations are detailed in Appendix B.
 4.2 Main Results
-Table 1 presents results evaluating CRAFT against open-source pre-trained models, closed-source API models, and SFT baselines across three multi-hop QA benchmarks.
-Additional experiments in Appendix C.
-CRAFT improves accuracy across models at all scales.
+Table 3.3.2 presents results across three multi-hop QA benchmarks.
+CRAFT improves accuracy across all model scales.
 For the 7B model, CRAFT achieves gains of +17.5%/+9.7%/+21.0% EM on MuSiQue/HotpotQA/2WikiMHQA compared to base Qwen2.5-7B, corresponding to relative improvements of 47%/17%/36%.
-The transformation is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model—a 16 improvement, showing how CRAFT’s decomposed rewards unlock latent reasoning capacity in small models.
-CRAFT improves faithfulness across models at all scales.
-For the 7B model, CRAFT achieves +22.0%/+8.9%/+16.1% EM improvements over SFT on MuSiQue/HotpotQA/2WikiMHQA.
-While SFT often degrades Faithfulness compared to base models (50.85% vs.
-58.25% for 7B on MuSiQue, a drop of -7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT.
-This divergence illustrates that SFT learns to imitate surface patterns, whereas CRAFT’s judge-based faithfulness reward encourages internally consistent, evidence-grounded traces.
-CRAFT matches or exceeds closed-source API models.
-54.35% EM on MuSiQue (vs.
+The improvement is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model, a 16 improvement, demonstrating that CRAFT’s decomposed rewards are effective even at very small scales.
+CRAFT improves faithfulness across all model scales.
+While SFT often degrades Faithfulness compared to base models (e.g., 50.85% vs.
+58.25% for 7B on MuSiQue, a drop of 7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT.
+CRAFT matches or exceeds larger closed-source API models.
+CRAFT achieves 54.35% EM on MuSiQue (vs.
 GPT-5-mini 53.45%), 66.61%/82.09% EM/F1 on HotpotQA (vs.
@@
 70.55%, +8.2%).
-This demonstrates that GRPO can train a 7B open-source model to match or exceed proprietary systems.
+This demonstrates that GRPO with decomposed rewards can train a 7B open-source model to match or exceed proprietary systems.
 4.3 Ablation Study
 We conduct two ablation studies: (1) trace template ablation with aligned deterministic rewards, and (2) judge reward ablation.
-In CRAFT, the trace template determines deterministic reward availability (, , ), so variants ablate both trace components and their corresponding rewards.
-Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>, disabling all deterministic rewards and judge reward.
-We further ablate the judge reward by training with and without .
-Table 2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
+Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>.
+Table 3.3.2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
 Removing structure degrades performance.
-For base models, answer-only CRAFT achieves the lowest EM (31.55%/53.40%/50.60% on MuSiQue/HotpotQA/2WikiMHQA), underperforming CRAFT (38.35%/59.30%/60.80%) by 5.9%–10.2%.
-While GRPO improves CRAFT substantially to 56.25%/63.25%/67.15%, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%).
-Notably, CRAFT cannot be evaluated for Faithfulness as it lacks reasoning traces.
-Trace components reveal accuracy-faithfulness trade-offs.
-For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without explicit training.
-After GRPO training, this gap narrows substantially (CRAFT: 80.35%/95.00%/94.80% vs.
-CRAFT: 86.70%/96.00%/94.10%), while CRAFTv1/CRAFTv2 improve auditability via explicit planning and citation.
-CRAFT outperforms SFT across all configurations.
-CRAFT achieves gains across all variants: +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models.
-In contrast, SFT shows frequent degradation (up to -6.6% EM on CRAFT) and systematic Faithfulness drops (-7.4% to -10.0% on MuSiQue).
-For CRAFT, GRPO improves MuSiQue EM by +17.5% (36.85%54.35%) while SFT decreases it by -4.5% (36.85%32.35%).
+Answer-only CRAFT achieves the lowest base EM (31.55%/53.40%/50.60%), underperforming CRAFT by 5.9%–10.2%.
+While GRPO improves CRAFT substantially, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%).
+Simpler traces achieve higher base faithfulness but lower auditability.
+For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without training.
+After GRPO, this gap narrows substantially, while CRAFT/CRAFT improve auditability via explicit planning and citation.
+SFT degrades faithfulness despite improving format compliance.
+CRAFT achieves +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models, while SFT shows frequent degradation (up to 6.6% EM, 10.0% Faithfulness on MuSiQue).
 Judge reward is essential for faithfulness.
-Without judge reward, Faithfulness drops -8.9% to -9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable.
-This validates our dual reward design: deterministic rewards ensure structural correctness and answer accuracy, while judge rewards are essential for semantic faithfulness.
-4.4 Training and Traces Analysis
-| Question: Which member of the research project conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005 was Kevin R. Henke most critical of? |
-| Gold Answer: Russell Humphreys |
-| plan |
-| Subquestion1: Which research project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005? |
-| Subquestion2: Who was Kevin R. Henke most critical of in [ANSWER_0]? |
-| gold_docs |
-| [2, 7] |
-| reason |
-| Subquestion1: From document [2], it is clear that the RATE project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005. |
-| Subquestion2: Document [7] states that Kevin R.
-Henke has been critical of the RATE project’s results, specifically mentioning that he has argued that Russell Humphreys, a young-earth creationist involved in the project, has made errors in his research.
-|
-| answer |
-| Russell Humphreys ✓ |
-4.4.1 Training Dynamics
+Without judge reward, Faithfulness drops 8.9% to 9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable.
+This validates our dual reward design: deterministic rewards ensure structural correctness, while judge rewards are essential for semantic faithfulness.
+Notably, adding the judge reward also improves EM (e.g., +3.0% on CRAFT MuSiQue), confirming that faithful reasoning supervision enhances rather than compromises answer accuracy.
+4.4 Training Dynamics
+Figure 3 shows the training dynamics of GRPO across model scales using CRAFT(full format).
 Training performance scales with model size.
-Figure 3 shows total reward rising monotonically with scale: 0.5B (), 1.5B (), 3B (), and 7B ().
-Across all scales, reward curves converge smoothly without oscillation, indicating stable GRPO optimization under our multi-reward design.
-Notably, the largest relative gain appears at 1.5B, while marginal improvements diminish at 7B, suggesting saturation once most reward components approach their upper bounds.
-Sub-rewards reveal fast format learning and a faithfulness threshold.
-Format Reward () converges within 50 training steps across all scales (0.5B: ; 7B remains near throughout), confirming that structural constraints are quickly internalized once the output template is learned.
-Answer Reward () scales with model capacity, reaching final scores of 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B, while Gold_doc Reward () improves steadily (0.050.27; 0.120.66; 0.190.55; 0.760.92), reflecting stronger evidence grounding in larger models.
-Faithfulness exhibits a clear capacity threshold.
-Faithfulness Reward () remains near zero for 0.5B () and reaches only modest levels at 1.5B (0.10), whereas 3B and 7B show sharp improvements (, ).
-This transition indicates that structured consistency auditing is substantially more demanding than surface format compliance or answer correctness, and only emerges once the model has sufficient capacity to internalize and satisfy judge-based semantic constraints.
-4.4.2 Output Traces Case Study
-We analyze the quality of generated reasoning traces to understand how CRAFT structures multi-hop reasoning.
-Table 3 presents a representative example using the CRAFT, demonstrating CRAFT’s systematic approach with explicit planning, gold document citation, step-by-step reasoning, and final answer extraction.
-Additional case study in Appendix A.
-About traces format analysis in Appendix C.3.
+Total reward increases monotonically with scale, and all curves converge smoothly, indicating stable optimization under our multi-reward design.
+Larger models achieve both higher starting points and higher ceilings, with the 7B model reaching a total reward of 3.41 compared to 1.47 for 0.5B.
+Format compliance is learned quickly; faithfulness requires sufficient model capacity.
+converges within 50 steps across all scales, indicating that structural constraints are the easiest to satisfy.
+scales with capacity, reaching 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B; notably, 1.5B and 3B achieve similar answer accuracy but diverge sharply on faithfulness.
+remains near zero for 0.5B and reaches only modest levels at 1.5B, whereas 3B and 7B show clear improvements (, ).
+This indicates that faithfulness learning is substantially more demanding than format compliance or answer accuracy, and only emerges with sufficient model capacity.
+CRAFT achieves near-perfect format compliance across all scales.
+Full format compliance results across all variants are reported in Appendix B.4.
+| Aspect | Both | Both | LLM | LLM | Agr. | |
+| Pass | Fail | Human | Human | |||
+| Plan-Reason | 393 | 78 | 15 | 14 | 94.2% | 0.81 |
+| Evidence Gr. | 375 | 83 | 25 | 17 | 91.6% | 0.75 |
+| Answer Deriv. | 410 | 71 | 9 | 10 | 96.2% | 0.86 |
+| Gold Doc Cit. | 370 | 80 | 33 | 17 | 90.0% | 0.70 |
+| Overall (Avg.) | 387 | 78 | 20 | 15 | 93.0% | 0.78 |
+4.5 Human Validation of Judge Reliability
+To validate our LLM-based faithfulness metric, two annotators with NLP research experience independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT using the same four-aspect binary criteria as the LLM judge (Algorithm 1).
+We select MuSiQue as it is the most challenging benchmark, requiring up to four reasoning hops.
+Results show 93.0% overall agreement with Cohen’s Artstein (2017), indicating substantial agreement (Table 3).
+Answer Derivation achieves the highest agreement (), while Gold Document Citation shows lower agreement (), reflecting inherent ambiguity in whether cited documents provide essential versus supplementary support.
+These results confirm that the LLM judge aligns well with human judgment in absolute terms.
+We further re-evaluate all configurations with an independent judge (Gemini-2.5-Flash) in Appendix C; model rankings are perfectly preserved (Spearman = 1.0).
+Together, these two evaluations address the circularity of using Qwen3-Max for both training and evaluation: human validation confirms that the judge’s criteria reflect genuine faithfulness rather than idiosyncratic preferences, while cross-judge consistency confirms that the reported improvements are not artifacts of a specific evaluator.
+| Aspect | Base (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
+[…]
+Subq1: Who is the creator?
+Subq2: What is the nationality? Subq3: What is the avg income? |
+Subq1: Who is the creator?
+Subq2: What is the avg income for same nationality? |
+[…]
+[…]
+[…]
+[…]
+| 0 - Concludes “English” despite identifying Spelling | 1 - creatornationalityincome | |
+| 0 - Misuses doc [5] (Crosby creator) | 1 - All cited docs in gold_docs | |
+| 0 - Derives “English” but uses US income | 1 - $72,641 supported by reasoning | |
+| 0 - “Creator is English” unsupported | 1 - All claims verified | |
+| 0.0 | 1.0 | |
+4.6 Case Study
+Table 4 presents a case from MuSiQue where the base model produces the correct answer through unfaithful reasoning (EM=1, =0), while CRAFT achieves both correctness and faithfulness (EM=1, =1).
+The base model answers “$72,641” correctly but confuses Aaron Spelling with Lorraine Crosby from a distractor document, concluding the creator is “English” yet using US income data.
+This is a typical distractor-induced entity confusion: the model latches onto a surface-similar passage and constructs a plausible-sounding but internally inconsistent chain.
+Notably, all four judge checks fail (====0), demonstrating how the chain of faithfulness catches errors that outcome-level metrics miss entirely.
+In contrast, CRAFT correctly traces SpellingAmericanUS income, citing only relevant documents and passing all four faithfulness checks.
+An additional case study where both accuracy and faithfulness are corrected is provided in Appendix D.
 5 Conclusion
 We present CRAFT, a GRPO-based reinforcement learning framework that trains RAG-based LLMs to generate structured, machine-auditable reasoning traces for multi-hop QA.
-CRAFT employs a dual reward design that combines deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness.
-Experiments on three benchmarks show that CRAFT consistently improves both accuracy and faithfulness over base models and SFT, with CRAFT matching or exceeding closed-source API models.
-Ablation results highlight the importance of structured traces and judge rewards for learning faithful reasoning, while training dynamics analysis indicates that faithfulness learning requires sufficient model capacity.
-Appendix A Faithfulness Case Study
-We present two representative cases from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs, demonstrating how CRAFT training improves both answer correctness and reasoning faithfulness.
-Each case includes the complete judge evaluation across all four faithfulness dimensions defined in Algorithm 1.
-A.1 Case 1: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
-This case demonstrates how an unfaithful and incorrect response is transformed into a faithful and correct one.
-Question: Who won the Indy car race in the largest city of the state where the Oh Yeah album performer was born?
-Gold Answer: Mario Andretti
-| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
-| Plan |
-Subq1: Where is the Oh Yeah album from?
-Subq2: What is the largest city and capital of that state? Subq3: Who won the Indy car race in that city? |
-Subq1: Where is the Oh Yeah performer from?
-Subq2: What is the largest city and capital of the state where the Oh Yeah performer is from? Subq3: Who won the Indy car race in the largest city and capital of that state? |
-| Gold Docs | [5, 12, 14] | [11, 12, 1] |
-| Reason | [Answer1] Charles Mingus, the creator of the Oh Yeah album, was born in Nogales, Arizona (doc [5]).
-[Answer2] The largest city and capital of Arizona is Phoenix (doc [12]).
-[Answer3] The Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and the winner of the Indy car race there in 2016 was Ryan Hunter-Reay (doc [1]).
-| Doc [11] indicates that the Oh Yeah performer, Charles Mingus, is from Nogales, Arizona.
-Doc [12] states that the capital and largest city of Arizona is Phoenix.
-Doc [1] mentions that the Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and it is an Indy car race where Mario Andretti won.
-|
-| Answer | Ryan Hunter-Reay | Mario Andretti ✓ |
-| Judge Evaluation | ||
-| Plan-Reason () | 1 - Plan steps followed in order | 1 - Plan followed: performer→state→city→winner |
-| Receipt-Reason () | 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs [11,12,1] in gold_docs |
-| Reason-Answer () | 1 - Answer matches reasoning conclusion | 1 - “Mario Andretti” supported by reasoning |
-| Evidence Grounded () | 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified in cited docs |
-| Overall () | 0.0 - Failures: doc_outside_gold_docs, evidence_not_supported | 1.0 - All checks passed |
-A.2 Case 2: Base (EM=1, Faith=0) CRAFT (EM=1, Faith=1)
-This case demonstrates that correct answers can arise from unfaithful reasoning, and CRAFT enforces faithfulness even when accuracy is preserved.
-Question: What is the average yearly income for a person of the same nationality as the creator of MacGruder and Loud?
-Gold Answer: $72,641
-| Aspect | Base Model (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
-| Plan |
-Subq1: Who is the creator of MacGruder and Loud?
-Subq2: What is the nationality of [Answer1]? Subq3: What is the average yearly income for a person of [Answer2] nationality? |
-Subq1: Who is the creator of MacGruder and Loud?
-Subq2: What is the average yearly income for a person of the same nationality as the creator? |
-| Gold Docs | [4, 5, 2] | [8, 4] |
-| Reason | [Answer1]: From document 4, we learn that the creator of MacGruder and Loud is Aaron Spelling.
-Document 5 tells us that Lorraine Crosby, known as “Mrs.
-Loud,” is English.
-Therefore, the nationality of the creator of MacGruder and Loud is English.
-Document 2 provides the average yearly income for Americans, which is the closest nationality we have data for.
-The mean household income in the United States was $72,641.
-| Doc [8] states that MacGruder and Loud was one of the few failures from Aaron Spelling’s production company.
-Doc [4] provides information about Aaron Spelling, indicating that he was an American film and television producer.
-Therefore, the creator of MacGruder and Loud is American.
-The average yearly income for an American, as mentioned in doc [2], is $72,641.
-|
-| Answer | $72,641 ✓ (correct answer, but unfaithful reasoning) | $72,641 ✓ |
-| Judge Evaluation | ||
-| Plan-Reason () | 0 - Identifies Aaron Spelling but concludes “English” | 1 - Plan followed: creator→nationality→income |
-| Receipt-Reason () | 0 - Misuses doc [5] (Lorraine Crosby creator) | 1 - All cited docs [8,4] in gold_docs |
-| Reason-Answer () | 0 - Derives “English” but uses US income | 1 - “$72,641” supported by reasoning |
-| Evidence Grounded () | 0 - “Creator is English” unsupported | 1 - Spelling→American→US income verified |
-| Overall () | 0.0 - Failures: plan_ignored, answer_not_supported, evidence_not_supported | 1.0 - All checks passed |
-A.3 Summary
-These cases illustrate two critical aspects of CRAFT’s faithfulness enforcement:
-Case 1 shows that CRAFT corrects both unfaithful reasoning (citing documents outside gold_docs, hallucinating unsupported claims) and incorrect answers.
-Case 2 demonstrates that answer correctness does not imply reasoning faithfulness.
-The Base model produces the correct answer through a flawed reasoning chain (confusing Lorraine Crosby with Aaron Spelling), while CRAFT ensures the answer is derived through a consistent, evidence-grounded process.
-The judge evaluation tables reveal that CRAFT training optimizes all four faithfulness dimensions simultaneously: plan-reason consistency (), receipt-reason consistency (), reason-answer consistency (), and evidence grounding ().
-This multi-dimensional supervision prevents shortcut learning and ensures that models produce not just correct answers, but trustworthy reasoning traces.
-Appendix B Implementation Details
+By combining deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness, CRAFT consistently improves both accuracy and faithfulness across model scales, with CRAFT matching or exceeding closed-source API models.
+Notably, semantic judge-based rewards improve answer accuracy rather than compromise it.
+Limitations
+Our human validation of judge reliability is conducted on 500 samples from MuSiQue only.
+While we select MuSiQue as the most challenging benchmark and supplement with cross-judge evaluation (Appendix C), extending human annotation to all three datasets would further strengthen the reliability analysis; we leave this to future work due to the cost of expert annotation on multi-hop reasoning traces.
+Second, due to the substantial computational cost of GRPO training with online judge inference, our experiments are limited to models up to 7B parameters trained on 20K samples from three multi-hop QA benchmarks (HotpotQA, 2WikiMHQA, MuSiQue).
+A central finding of this work is that faithfulness supervision can improve rather than compromise answer accuracy.
+We focus on establishing this finding within the training distribution first, as confirming that the accuracy–faithfulness synergy holds under controlled conditions is a necessary prerequisite before testing its generalization.
+Scaling to larger models and evaluating whether this synergy transfers to out-of-distribution QA domains remains important future work.
+Appendix A Implementation Details
+This section expands on the experimental setup described in §4.
 Training Framework.
 We implement CRAFT using MS-Swift Zhao et al.
-[2025], an efficient and scalable fine-tuning framework for large language models.
+(2025), an efficient fine-tuning framework for large language models.
 MS-Swift provides native support for GRPO training with vLLM acceleration, enabling efficient on-policy sampling during reinforcement learning.
-All experiments are conducted on NVIDIA H100 96GB x4 GPUs.
+All experiments are conducted on NVIDIA H100 96GB 4 GPUs.
+Table A summarizes the key hyperparameters.
 Model Configuration.
@@
 All models use bfloat16 precision with Flash Attention 2 for memory-efficient training.
-Table 6 summarizes the key hyperparameters.
+For SFT baselines, we fine-tune each model on the same training data using standard cross-entropy loss with a learning rate of 2e-5 and 3 epochs.
 | Hyperparameter | Value |
-| GRPO Configuration | |
+[…]
 | Batch Generations Accum. | 4 4 64 |
@@
 | Max Completion Length | 256 tokens |
-| Optimization | |
+[…]
 | Learning Rate | 5e-6 |
@@
 | Warmup Ratio | 0.1 |
-| Infrastructure | |
+[…]
 | Precision | bfloat16 |
@@
 Judge Model.
-For the faithfulness reward and evaluation, we employ Qwen3-Max Team [2025] as the LLM judge.
-This model provides reliable semantic consistency assessment while maintaining computational efficiency.
-Because Qwen3-Max already functions as the judge for reward computation and evaluation, it is excluded from all performance comparison experiments to avoid role overlap.
+We employ Qwen3-Max Team (2025) as the LLM judge for both the faithfulness reward during training and the Faithfulness metric during evaluation.
+The judge is prompted with the query, retrieved documents, and the model’s trace, and returns binary scores for each applicable audit dimension.
+We use temperature 0.0 for deterministic judge outputs.
+Because Qwen3-Max serves dual roles as reward provider and evaluator, it is excluded from all performance comparisons.
+We verify that this dual usage does not bias our conclusions through cross-judge evaluation with Gemini-2.5-Flash (Appendix C).
 Reward Function Configuration.
-The reward functions vary across prompt templates based on their structural components:
-CRAFT/CRAFT: , , (includes <gold_docs> field)
-CRAFT/CRAFT: , (no explicit document citation)
-CRAFT: , (direct answer generation)
-All reward weights are set to 1.0, treating each component equally.
-The faithfulness reward is computed via LLM judge during evaluation but not used as a training signal in the ablation experiments (Table 2, “w/o judge reward”).
+The active reward components vary across trace templates based on the available structural fields: CRAFT/CRAFT: , , , (full supervision); CRAFT/CRAFT: , , (no citation reward as <gold_docs> is absent); CRAFT: , only (no reasoning fields for faithfulness audit).
+All reward weights are set to 1.0.
+This is a natural default since , , , and all produce values in ( is binary, and are F1 scores, and is an average of binary checks), so equal weighting does not introduce scale imbalance.
+Moreover, converges to 1.0 within 50 training steps across all scales (Figure 3), contributing negligible gradient signal thereafter.
+In the ablation without judge reward (Table 3.3.2, “w/o judge reward”), is removed from training but still computed during evaluation.
 Training Data.
-Each prompt template uses a dedicated training dataset of 20,000 samples, formatted according to the template’s structural requirements.
-The data distribution is: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA samples.
-Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns that indicate diverse linguistic characteristics across the three multi-hop QA datasets.
-Appendix C Additional Experiments
-This section presents comprehensive results for prompt template variants CRAFT–CRAFT, including training dynamics and evaluation results across all model scales.
-C.1 Training Dynamics
-Figures 5–8 illustrate the training dynamics for CRAFT–CRAFT templates.
-Consistent with CRAFT (main paper Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
-Template-specific patterns: (1) CRAFT (Figure 5) shows similar dynamics to CRAFT, with all four reward components (, , , ) contributing to optimization.
-(2) CRAFT/CRAFT (Figures 6–7) lack but maintain stable training through format and answer rewards.
-(3) CRAFT (Figure 8) uses only and —notably, the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
-Faithfulness reward verification.
-To verify that the low training faithfulness reward (Figure 3) reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics (Algorithm 1) on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0).
-Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%.
-This indicates that while 69.7% of correct predictions exhibit fully faithful reasoning, 30.3% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness.
-This test-time value is approximately six times higher than the training-time average of 0.11 observed for this model (Figure 3), confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, resulting in lower average rewards during training, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
+Each prompt template uses a dedicated training dataset of 20,000 samples: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA.
+We oversample MuSiQue as it contains questions requiring 2–4 reasoning hops, providing greater diversity in reasoning depth compared to the predominantly 2-hop questions in HotpotQA and 2WikiMHQA.
+Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns across the three datasets.
 | Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
-| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
-| Open-source | ||||||||||
-| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
-| Qwen2.5-1.5B Team [2024] | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
-| Qwen2.5-3B Team [2024] | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
-| Qwen2.5-7B Team [2024] | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
-| Qwen3-4B Team [2025] | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
-| Qwen3-30B Team [2025] | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
-| API | ||||||||||
-| GPT-5-mini OpenAI [2025] | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
-| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
-| Gemini-2.5-Flash Google DeepMind [2025] | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
-| SFT | ||||||||||
-| Qwen2.5-0.5B Team [2024] | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
-| Qwen2.5-1.5B Team [2024] | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
-| Qwen2.5-3B Team [2024] | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
-| Qwen2.5-7B Team [2024] | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
-| Ours (RL) | ||||||||||
+| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
+[…]
+| Qwen2.5-0.5B | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
+| Qwen2.5-1.5B | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
+| Qwen2.5-3B | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
+| Qwen2.5-7B | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
+| Qwen3-4B | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
+| Qwen3-30B | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
+[…]
+| GPT-5-mini | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
+| Deepseek-V3.2 | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
+| Gemini-2.5-Flash | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
+[…]
+| Qwen2.5-0.5B | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
+| Qwen2.5-1.5B | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
+| Qwen2.5-3B | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
+| Qwen2.5-7B | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
+[…]
 | CRAFT | Trained | 6.50 | 12.76 | 9.85 | 20.60 | 29.86 | 23.70 | 22.65 | 28.79 | 15.90 |
@@
 | Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
-| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
-| Open-source | ||||||||||
-| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
-| Qwen2.5-1.5B Team [2024] | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
-| Qwen2.5-3B Team [2024] | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
-| Qwen2.5-7B Team [2024] | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
-| Qwen3-4B Team [2025] | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
-| Qwen3-30B Team [2025] | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
-| API | ||||||||||
-| GPT-5-mini OpenAI [2025] | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
-| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
-| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
-| SFT | ||||||||||
-| Qwen2.5-0.5B Team [2024] | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
-| Qwen2.5-1.5B Team [2024] | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
-| Qwen2.5-3B Team [2024] | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
-| Qwen2.5-7B Team [2024] | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
-| Ours (RL) | ||||||||||
+| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
+[…]
+| Qwen2.5-0.5B | Pre-Trained | 0.25 | 1.72 | 0.85 | 1.80 | 7.31 | 6.40 | 1.85 | 8.81 | 4.60 |
+| Qwen2.5-1.5B | Pre-Trained | 2.15 | 6.98 | 15.55 | 7.30 | 16.58 | 35.20 | 10.05 | 17.10 | 21.35 |
+| Qwen2.5-3B | Pre-Trained | 17.70 | 26.19 | 38.85 | 38.85 | 51.77 | 69.70 | 34.30 | 43.93 | 62.00 |
+| Qwen2.5-7B | Pre-Trained | 38.95 | 49.51 | 68.70 | 57.75 | 72.10 | 90.65 | 60.65 | 69.82 | 84.25 |
+| Qwen3-4B | Pre-Trained | 29.75 | 39.04 | 57.40 | 54.60 | 70.04 | 88.60 | 56.20 | 66.71 | 86.80 |
+| Qwen3-30B | Pre-Trained | 43.80 | 53.59 | 72.65 | 61.05 | 75.94 | 93.80 | 63.95 | 73.70 | 90.30 |
+[…]
+| GPT-5-mini | API | 52.45 | 67.83 | 89.45 | 60.98 | 79.77 | 98.20 | 69.60 | 80.28 | 97.25 |
+| Deepseek-V3.2 | API | 52.75 | 67.04 | 83.40 | 62.90 | 80.00 | 97.10 | 68.90 | 78.06 | 97.65 |
+| Gemini-2.5-Flash | API | 51.95 | 63.53 | 82.65 | 64.00 | 79.59 | 96.00 | 70.15 | 79.45 | 96.85 |
+[…]
+| Qwen2.5-0.5B | Trained | 1.10 | 5.92 | 5.35 | 12.80 | 21.39 | 17.45 | 22.45 | 27.18 | 17.95 |
+| Qwen2.5-1.5B | Trained | 9.65 | 17.17 | 15.55 | 36.15 | 47.65 | 45.10 | 36.00 | 43.19 | 38.00 |
+| Qwen2.5-3B | Trained | 22.15 | 31.93 | 43.85 | 51.40 | 65.75 | 79.90 | 53.40 | 61.93 | 71.90 |
+| Qwen2.5-7B | Trained | 35.60 | 47.66 | 61.35 | 59.25 | 73.68 | 88.25 | 65.40 | 74.33 | 85.80 |
+[…]
 | CRAFT | Trained | 4.75 | 10.09 | 7.95 | 20.20 | 30.14 | 20.85 | 20.50 | 25.65 | 12.50 |
@@
 | Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
-| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
-| Open-source | ||||||||||
-| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
-| Qwen2.5-1.5B Team [2024] | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
-| Qwen2.5-3B Team [2024] | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
-| Qwen2.5-7B Team [2024] | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
-| Qwen3-4B Team [2025] | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
-| Qwen3-30B Team [2025] | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
-| API | ||||||||||
-| GPT-5-mini OpenAI [2025] | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
-| Deepseek-V3.2 DeepSeek-AI [2025] | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
-| Gemini-2.5-Flash Google DeepMind [2025] | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
-| SFT | ||||||||||
-| Qwen2.5-0.5B Team [2024] | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
-| Qwen2.5-1.5B Team [2024] | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
-| Qwen2.5-3B Team [2024] | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
-| Qwen2.5-7B Team [2024] | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
-| Ours (RL) | ||||||||||
+| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
+[…]
+| Qwen2.5-0.5B | Pre-Trained | 0.35 | 2.69 | 4.85 | 3.00 | 7.51 | 14.10 | 5.05 | 9.98 | 8.00 |
+| Qwen2.5-1.5B | Pre-Trained | 4.85 | 9.78 | 16.65 | 26.80 | 36.98 | 44.65 | 21.65 | 28.33 | 31.15 |
+| Qwen2.5-3B | Pre-Trained | 21.95 | 32.30 | 41.15 | 48.30 | 61.44 | 76.05 | 41.95 | 50.67 | 60.80 |
+| Qwen2.5-7B | Pre-Trained | 38.35 | 49.29 | 76.60 | 59.30 | 74.40 | 93.75 | 60.80 | 70.20 | 86.15 |
+| Qwen3-4B | Pre-Trained | 38.50 | 47.61 | 77.55 | 60.70 | 76.06 | 95.60 | 65.40 | 74.61 | 93.90 |
+| Qwen3-30B | Pre-Trained | 46.75 | 56.34 | 84.95 | 63.35 | 78.13 | 97.10 | 70.35 | 79.04 | 95.95 |
+[…]
+| GPT-5-mini | API | 52.05 | 68.10 | 93.15 | 60.85 | 79.83 | 98.05 | 70.05 | 81.09 | 97.10 |
+| Deepseek-V3.2 | API | 53.20 | 67.96 | 88.25 | 61.55 | 78.67 | 98.25 | 70.35 | 79.35 | 98.55 |
+| Gemini-2.5-Flash | API | 51.25 | 62.92 | 90.20 | 63.75 | 79.65 | 97.20 | 70.60 | 79.41 | 97.80 |
+[…]
+| Qwen2.5-0.5B | Trained | 0.50 | 4.62 | 13.50 | 11.75 | 21.27 | 37.40 | 23.15 | 28.68 | 26.85 |
+| Qwen2.5-1.5B | Trained | 7.25 | 12.85 | 16.40 | 33.25 | 44.48 | 51.95 | 36.30 | 43.53 | 42.70 |
+| Qwen2.5-3B | Trained | 23.60 | 33.10 | 50.15 | 50.25 | 64.37 | 83.85 | 55.90 | 64.77 | 74.10 |
+| Qwen2.5-7B | Trained | 31.80 | 42.71 | 66.60 | 57.90 | 72.43 | 89.35 | 63.15 | 72.00 | 82.25 |
+[…]
 | CRAFT | Trained | 8.45 | 15.11 | 25.40 | 25.95 | 36.20 | 40.70 | 24.70 | 30.94 | 26.50 |
@@
 | EM | F1 | EM | F1 | EM | F1 | ||
-| Open-source | |||||||
-| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
-| Qwen2.5-1.5B Team [2024] | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
-| Qwen2.5-3B Team [2024] | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
-| Qwen2.5-7B Team [2024] | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
-| Qwen3-4B Team [2025] | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
-| Qwen3-30B Team [2025] | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
-| API | |||||||
-| GPT-5-mini OpenAI [2025] | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
-| Deepseek-V3.2 DeepSeek-AI [2025] | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
-| Gemini-2.5-Flash Google DeepMind [2025] | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
-| SFT | |||||||
-| Qwen2.5-0.5B Team [2024] | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
-| Qwen2.5-1.5B Team [2024] | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
-| Qwen2.5-3B Team [2024] | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
-| Qwen2.5-7B Team [2024] | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
-| Ours (RL) | |||||||
+[…]
+| Qwen2.5-0.5B | Pre-Trained | 0.30 | 3.20 | 5.45 | 12.18 | 6.40 | 12.37 |
+| Qwen2.5-1.5B | Pre-Trained | 10.20 | 17.58 | 35.95 | 47.77 | 26.05 | 33.29 |
+| Qwen2.5-3B | Pre-Trained | 23.20 | 33.31 | 49.30 | 62.58 | 41.45 | 50.09 |
+| Qwen2.5-7B | Pre-Trained | 31.55 | 41.97 | 53.40 | 67.92 | 50.60 | 59.29 |
+| Qwen3-4B | Pre-Trained | 39.55 | 49.23 | 60.40 | 75.32 | 57.70 | 67.43 |
+| Qwen3-30B | Pre-Trained | 49.30 | 59.33 | 63.05 | 78.33 | 64.70 | 73.71 |
+[…]
+| GPT-5-mini | API | 51.10 | 64.43 | 63.63 | 80.19 | 65.28 | 76.53 |
+| Deepseek-V3.2 | API | 58.10 | 70.97 | 64.15 | 79.91 | 67.45 | 76.66 |
+| Gemini-2.5-Flash | API | 54.80 | 66.64 | 65.68 | 80.79 | 71.85 | 80.59 |
+[…]
+| Qwen2.5-0.5B | Trained | 0.15 | 0.85 | 0.80 | 2.85 | 2.80 | 4.70 |
+| Qwen2.5-1.5B | Trained | 10.05 | 17.23 | 34.10 | 45.92 | 34.30 | 40.92 |
+| Qwen2.5-3B | Trained | 19.60 | 28.60 | 46.60 | 60.23 | 50.40 | 57.75 |
+| Qwen2.5-7B | Trained | 27.00 | 37.82 | 53.30 | 68.12 | 56.15 | 63.79 |
+[…]
 | CRAFT | Trained | 0.00 | 0.00 | 0.00 | 0.00 | 0.05 | 0.05 |
@@
 | Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | Base | GRPO | ||
-| Qwen2.5-0.5B | |||||||||||
+[…]
 | Avg. | 0.0 | 98.3(+98.3) | 1.1 | 99.2(+98.1) | 0.0 | 94.3(+94.3) | 1.0 | 99.4(+98.4) | 9.0 | 0.1(-8.9) | |
-| HotpotQA | 0.0 | 99.6(+99.6) | 0.1 | 99.8(+99.7) | 0.1 | 97.5(+97.4) | 0.8 | 99.8(+99.0) | 5.7 | 0.1(-5.6) | |
-| 2WikiMQA | 0.0 | 99.5(+99.5) | 3.2 | 99.7(+96.5) | 0.0 | 92.5(+92.5) | 2.3 | 99.8(+97.5) | 16.5 | 0.1(-16.4) | |
-| MuSiQue | 0.0 | 95.8(+95.8) | 0.0 | 98.2(+98.2) | 0.0 | 92.8(+92.8) | 0.0 | 98.5(+98.5) | 4.9 | 0.0(-4.9) | |
-| Qwen2.5-1.5B | |||||||||||
+| HotpotQA | 0.0 | 99.6 | 0.1 | 99.8 | 0.1 | 97.5 | 0.8 | 99.8 | 5.7 | 0.1 | |
+| 2WikiMQA | 0.0 | 99.5 | 3.2 | 99.7 | 0.0 | 92.5 | 2.3 | 99.8 | 16.5 | 0.1 | |
+| MuSiQue | 0.0 | 95.8 | 0.0 | 98.2 | 0.0 | 92.8 | 0.0 | 98.5 | 4.9 | 0.0 | |
+[…]
 | Avg. | 0.0 | 99.3(+99.3) | 2.1 | 95.1(+93.0) | 0.2 | 99.5(+99.3) | 1.2 | 81.9(+80.7) | 0.8 | 100.0(+99.2) | |
-| HotpotQA | 0.0 | 99.5(+99.5) | 2.9 | 98.9(+96.0) | 0.1 | 99.2(+99.1) | 1.0 | 92.9(+91.9) | 0.7 | 100.0(+99.3) | |
-| 2WikiMQA | 0.0 | 99.4(+99.4) | 0.8 | 99.2(+98.4) | 0.0 | 99.5(+99.5) | 1.4 | 91.2(+89.8) | 1.4 | 100.0(+98.6) | |
-| MuSiQue | 0.1 | 99.1(+99.0) | 2.5 | 87.4(+84.9) | 0.6 | 99.7(+99.1) | 1.2 | 61.6(+60.4) | 0.2 | 99.9(+99.7) | |
-| Qwen2.5-3B | |||||||||||
+| HotpotQA | 0.0 | 99.5 | 2.9 | 98.9 | 0.1 | 99.2 | 1.0 | 92.9 | 0.7 | 100.0 | |
+| 2WikiMQA | 0.0 | 99.4 | 0.8 | 99.2 | 0.0 | 99.5 | 1.4 | 91.2 | 1.4 | 100.0 | |
+| MuSiQue | 0.1 | 99.1 | 2.5 | 87.4 | 0.6 | 99.7 | 1.2 | 61.6 | 0.2 | 99.9 | |
+[…]
 | Avg. | 13.2 | 96.9(+83.7) | 50.3 | 98.8(+48.5) | 6.4 | 98.0(+91.6) | 15.6 | 98.1(+82.5) | 41.6 | 100.0(+58.4) | |
-| HotpotQA | 14.8 | 98.2(+83.4) | 55.2 | 99.2(+44.0) | 4.3 | 99.4(+95.1) | 27.6 | 99.2(+71.6) | 32.8 | 100.0(+67.2) | |
-| 2WikiMQA | 7.7 | 99.0(+91.3) | 33.8 | 99.4(+65.6) | 6.3 | 99.1(+92.8) | 8.2 | 98.8(+90.6) | 23.9 | 100.0(+76.1) | |
-| MuSiQue | 17.2 | 93.5(+76.3) | 61.8 | 97.6(+35.8) | 8.7 | 95.6(+86.9) | 10.8 | 96.3(+85.5) | 68.1 | 100.0(+31.9) | |
-| Qwen2.5-7B | |||||||||||
+| HotpotQA | 14.8 | 98.2 | 55.2 | 99.2 | 4.3 | 99.4 | 27.6 | 99.2 | 32.8 | 100.0 | |
+| 2WikiMQA | 7.7 | 99.0 | 33.8 | 99.4 | 6.3 | 99.1 | 8.2 | 98.8 | 23.9 | 100.0 | |
+| MuSiQue | 17.2 | 93.5 | 61.8 | 97.6 | 8.7 | 95.6 | 10.8 | 96.3 | 68.1 | 100.0 | |
+[…]
 | Avg. | 94.1 | 99.3(+5.2) | 99.2 | 100.0(+0.8) | 94.4 | 97.2(+2.8) | 61.6 | 99.8(+38.2) | 15.4 | 100.0(+84.6) | |
-| HotpotQA | 96.1 | 100.0(+3.9) | 99.5 | 100.0(+0.5) | 95.7 | 99.1(+3.4) | 49.5 | 100.0(+50.5) | 11.2 | 100.0(+88.8) | |
-| 2WikiMQA | 95.4 | 99.1(+3.7) | 99.3 | 100.0(+0.7) | 96.0 | 97.0(+1.0) | 72.9 | 100.0(+27.1) | 9.6 | 100.0(+90.4) | |
-| MuSiQue | 90.9 | 98.9(+8.0) | 98.9 | 100.0(+1.1) | 91.6 | 95.5(+3.9) | 62.5 | 99.4(+36.9) | 25.4 | 100.0(+74.6) | |
-C.2 Evaluation Results
+| HotpotQA | 96.1 | 100.0 | 99.5 | 100.0 | 95.7 | 99.1 | 49.5 | 100.0 | 11.2 | 100.0 | |
+| 2WikiMQA | 95.4 | 99.1 | 99.3 | 100.0 | 96.0 | 97.0 | 72.9 | 100.0 | 9.6 | 100.0 | |
+| MuSiQue | 90.9 | 98.9 | 98.9 | 100.0 | 91.6 | 95.5 | 62.5 | 99.4 | 25.4 | 100.0 | |
+Appendix B Additional Results
+This section provides comprehensive supplementary results that expand on the main experiments in §4.
+We present training dynamics for all trace variants (§B.1), verification of faithfulness reward behavior (§B.2), full evaluation tables for CRAFT–CRAFT (§B.3), format compliance analysis (§B.4), format error patterns (§B.5), and model scale sensitivity (§B.6).
+B.1 Training Dynamics for Additional Variants
+Figures 5–8 show the training dynamics for CRAFT–CRAFT. Consistent with CRAFT (Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
+Template-specific patterns: (1) CRAFT shows similar dynamics to CRAFT, with all four reward components contributing.
+(2) CRAFT/CRAFT lack but maintain stable training.
+(3) CRAFT uses only and ; the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
+B.2 Faithfulness Reward Verification
+As noted in §4.4, the faithfulness reward remains low during training, particularly for smaller models.
+To verify that this reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0).
+Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%.
+This indicates that while 69.3% of correct predictions exhibit fully faithful reasoning, 30.7% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness.
+This test-time value is approximately six times higher than the training-time average of 0.11, confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
+B.3 Full Results for Additional Variants
+Tables A–A present the complete evaluation results for CRAFT–CRAFT across all model scales and baselines, complementing the CRAFT results in the main paper (Table 3.3.2).
 CRAFT Results.
-CRAFT retains the same four-field structure as CRAFT but with simplified instructions.
+CRAFT retains the same structure as CRAFT but removes <plan>.
 CRAFT achieves strong performance: 58.35% EM on MuSiQue (+24.0% vs Base), 67.28% on HotpotQA (+12.5% vs Base), and 79.70% on 2WikiMHQA (+23.7% vs Base).
@@
 CRAFT Results.
-CRAFT further simplifies the template while maintaining structured reasoning.
+CRAFT retains only <reason>+<answer>.
 CRAFT achieves the highest Faithfulness scores among all templates: 86.70% on MuSiQue, 96.00% on HotpotQA, and 94.10% on 2WikiMHQA.
-EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance when combined with appropriate reward design.
+EM scores (56.91%, 66.51%, 79.39%) remain competitive, demonstrating that simplified templates can achieve strong performance with appropriate reward design.
 CRAFT Results.
-CRAFT represents the minimal template with direct answer generation, lacking structured reasoning traces ( and not applicable).
+CRAFT uses direct answer generation without structured reasoning traces ( and not applicable).
 CRAFT achieves 56.25% EM on MuSiQue, 63.25% on HotpotQA, and 67.15% on 2WikiMHQA.
 A critical observation: CRAFT completely fails on this template (near-zero performance), indicating that smaller models require structured reasoning traces for stable training.
-This validates the importance of trace structure for model capacity constraints.
-CRAFT Template Sensitivity Analysis.
-Our experiments reveal a striking template sensitivity pattern unique to the smallest model, with different failure modes for SFT and GRPO training.
+B.4 Format Compliance Analysis
+Expanding on the format compliance summary in §4.4, Table A presents a comprehensive analysis across all CRAFT variants, model scales, and individual datasets.
+GRPO substantially improves format compliance across all scales.
+Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates, while even 3B models only reach 6.4%–50.3% on average.
+GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases.
+Even 7B base models benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%).
+This demonstrates that successfully teaches models to follow structured output requirements across all capacity levels.
+Smaller models struggle with complex templates but benefit most from GRPO training.
+For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a notable failure on CRAFT, where GRPO format compliance drops from 9.0% to 0.1%.
+We attribute this to a reward signal collapse: CRAFT uses only and , but the 0.5B model lacks sufficient capacity to produce correct answers without reasoning structure, yielding near-zero throughout training.
+With as the only effective reward, the model receives no learning signal to distinguish between the minimal <answer> format and the more complex XML structures encountered during exploration, leading to format drift rather than convergence.
+This failure is consistent with the complete accuracy collapse of CRAFT on CRAFT (Table A).
+In contrast, larger models master all formats: 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
+| Version | Base Model Errors (7B) | GRPO Model Errors (7B) |
+[…]
+Missing closing tags (60%):
+<answer>1958 Award missing </answer> Truncation (30%): response cut mid-sentence |
+Empty answer (rare):
+<answer> no content |
+[…]
+Missing closing tags (70%):
+<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): </answer> used for </reason> |
+No errors found ✓ |
+[…]
+Missing answer block (45%):
+entire <answer>...</answer> absent Truncation (40%): response cut off |
+Malformed tags (3%):
+truncated closing tag |
+[…]
+Missing closing tags (85%):
+<answer>Russell Humphreys missing </answer> |
+Extra text (0.2%):
+parenthetical note after tags |
+[…]
+Wrong format (85%):
+bare text without XML tags e.g., A Bug’s Life |
+No errors found ✓ |
+B.5 Format Error Analysis
+Table 11 complements the quantitative results in Table A by presenting representative error patterns across all CRAFT variants.
+Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
+B.6 CRAFT Template Sensitivity
+As observed in Table 3.3.2, CRAFT shows limited but non-trivial gains at the smallest scale.
+Here we examine this model’s sensitivity to trace template choice in detail, revealing striking differences between SFT and GRPO failure modes.
 SFT training collapses on templates CRAFT, CRAFT, and CRAFT (EM 1.5% across benchmarks), achieving meaningful performance only on CRAFT (1.10/12.80/22.45%) and CRAFT (0.50/11.75/23.15%).
@@
 Specifically, GRPO achieves 7.25/24.80/22.85% on CRAFT, 6.50/20.60/22.65% on CRAFT, 4.75/20.20/20.50% on CRAFT, and 8.45/25.95/24.70% on CRAFT, while CRAFT remains completely broken (0.00/0.00/0.05%).
-This dichotomy reveals two key insights: (1) SFT cold-start is fragile for 0.5B models—templates CRAFT/CRAFT with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse.
-(2) GRPO’s exploration mechanism overcomes SFT’s brittleness—reinforcement learning enables the model to discover valid reasoning paths even on complex structures (CRAFT/CRAFT), as long as some structural scaffolding exists.
-(3) Template CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure—direct answer generation lacks sufficient guidance for stable learning.
-The optimal template for 0.5B deployment is CRAFT (<plan>+<reason>+<answer> without <gold_docs>), which provides the best balance: enough structure to guide reasoning without exceeding representational capacity, and compatibility with both SFT initialization and GRPO refinement.
-C.3 Output Traces Format Analysis
-Table 11 presents a comprehensive analysis of format compliance across all CRAFT variants and model scales.
-GRPO dramatically improves format compliance across all scales.
-Base models struggle with structured output, particularly at smaller scales: 0.5B and 1.5B base models achieve near-zero format accuracy on complex templates (CRAFT–CRAFT), while even 3B models only reach 6.4%–50.3% on average.
-In stark contrast, GRPO training yields massive improvements: 0.5B models gain +92.5% to +99.7% on CRAFT–CRAFT templates, 1.5B models improve by +60.4% to +99.5%, and 3B models see +31.9% to +95.1% increases.
-Even 7B base models, despite already achieving 90.9%–99.5% on complex templates (CRAFT–CRAFT), benefit from GRPO with +0.5% to +8.0% gains, reaching near-perfect compliance (97.2%–100%).
-This demonstrates that reward successfully teaches models to follow structured output requirements across all capacity levels.
-Format complexity exhibits scale-dependent learnability.
-For simpler templates (CRAFT, CRAFT), smaller models show divergent patterns: 0.5B improves dramatically on CRAFT (+98.4%), but exhibits a puzzling anomaly on CRAFT, where GRPO models (0.1%) underperform base (9.0%) by -4.9% to -16.4% across datasets.
-This suggests the model defaults to more complex formats learned during multi-template training.
-In contrast, larger models master all formats: 1.5B achieves +80.7% to +99.3% gains, 3B reaches +58.4% to +91.6% improvements, and 7B models show the most dramatic gains on simple templates (+38.2% on CRAFT, +84.6% on CRAFT), indicating that base 7B models paradoxically struggle more with minimal-structure outputs than with complex multi-field templates.
-Format Error Analysis.
-Table 12 presents representative error patterns across all CRAFT variants.
-Base models predominantly fail due to missing closing tags (60%–85% of errors) and response truncation (30%–40% for complex templates), while GRPO dramatically reduces errors, with CRAFT and CRAFT achieving perfect compliance on 7B models.
-| Version | Base Model Errors (7B) | GRPO Model Errors (7B) |
-| CRAFT |
-Missing closing tags (60%):
-<answer>1958 Award missing </answer> Truncation (30%): ...grandmother of Henry St John ( response cut mid-sentence |
-Empty answer (rare):
-<answer> no content, missing </answer> |
-| CRAFT |
-Missing closing tags (70%):
-<answer>[Edwin Schlossberg] missing </answer> Tag mismatch (20%): <reason>...Michael Rubbo </answer> wrong closing tag None bare text |
-No errors found ✓ |
-| CRAFT |
-Missing <answer> block (45%):
-<reason>...answer is American. </reason> entire <answer>...</answer> missing Truncation (40%): ...cannot determine based on response cut off |
-Missing/malformed tags (3%):
-<answer>Infanta Maria missing </answer> <answer>The Woman in Fifth </ truncated closing tag |
-| CRAFT |
-Missing closing tags (85%):
-<answer>Russell Humphreys missing </answer> |
-Extra text after tags (0.2%):
-<answer>No</answer> (Note: The reasoning...) parenthetical note not allowed |
-| CRAFT |
-Wrong format (85%):
-A Bug’s Life no XML tags [Beaumont, Texas] bracket format [No] should be <answer>No</answer> |
-No errors found ✓ |
-C.4 Human Annotation Validation
-| Aspect | Both | Both | LLM | LLM | Agr. | |
-| Pass | Fail | Human | Human | |||
-| Plan-Reason | 388 | 83 | 18 | 11 | 94.2% | 0.81 |
-| Evidence Ground. | 380 | 80 | 23 | 17 | 92.0% | 0.75 |
-| Answer Deriv. | 400 | 80 | 10 | 10 | 96.0% | 0.86 |
-| Gold Doc Citation | 370 | 80 | 30 | 20 | 90.0% | 0.70 |
-| Overall | 385 | 80 | 20 | 15 | 93.0% | 0.78 |
-To validate the reliability of our LLM-based faithfulness metric, we conducted a human evaluation study following established protocols Adlakha et al.
-[2024]; Zheng et al.
-[2023].
-Two expert annotators independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT with CRAFT template, using the same four-aspect criteria as our LLM judge (Algorithm 1).
-We measure inter-annotator reliability using Cohen’s Kappa Artstein [2017]:
-where is the observed agreement and is the expected agreement by chance.
-Here denotes both-pass cases, denotes both-fail cases, (LLM pass, human fail) and (LLM fail, human pass) are disagreements, and is the total sample size.
-Table 13 presents the detailed per-aspect results.
-The results demonstrate consistently high agreement across all four aspects, with the overall faithfulness achieving 93.0% agreement and between automatic and human evaluation, indicating substantial agreement Zheng et al.
-[2023]; Bavaresco et al.
-[2024].
-Answer Derivation shows the highest agreement (), as the correctness of final answers is relatively unambiguous.
-Gold Document Citation has lower agreement (), reflecting occasional ambiguity in determining whether cited documents are strictly necessary.
-Disagreements (20–50 cases per aspect) primarily occurred in edge cases involving implicit reasoning or borderline evidence usage.
-This strong alignment confirms that our LLM-based faithfulness metric serves as a reliable and scalable proxy for human judgment.
+This dichotomy reveals three key insights: (1) SFT cold-start is fragile for 0.5B models: templates with explicit <gold_docs> fields overwhelm supervised imitation, causing training collapse.
+(2) GRPO’s exploration mechanism overcomes SFT’s brittleness: reinforcement learning enables the model to discover valid reasoning paths even on complex structures, as long as some structural scaffolding exists.
+(3) CRAFT’s complete failure across both training paradigms confirms that 0.5B models fundamentally require explicit reasoning structure for stable learning.
+The optimal template for 0.5B deployment is CRAFT (<reason>+<answer>), providing the best balance of structure and simplicity.
+| Model | Judge | MuSiQue | HotpotQA | 2WikiMHQA | |||
+| Faith. | Faith. | Faith. | |||||
+| Qwen2.5-7B | Qwen3-Max | 58.25 | – | 83.65 | – | 76.85 | – |
+| Gemini-Flash | 55.40 | -2.85 | 82.15 | -1.50 | 73.90 | -2.95 | |
+| Qwen2.5-7B | Qwen3-Max | 50.85 | – | 83.50 | – | 78.95 | – |
+| Gemini-Flash | 49.70 | -1.15 | 81.40 | -2.10 | 77.50 | -1.45 | |
+| CRAFT w/o judge | Qwen3-Max | 71.20 | – | 91.80 | – | 89.25 | – |
+| Gemini-Flash | 69.85 | -1.35 | 90.25 | -1.55 | 86.80 | -2.45 | |
+| CRAFT (full) | Qwen3-Max | 80.35 | – | 95.00 | – | 94.80 | – |
+| Gemini-Flash | 77.60 | -2.75 | 93.60 | -1.40 | 93.25 | -1.55 | |
+| Metric | MuSiQue | HotpotQA | 2WikiMHQA |
+| Pearson | 0.998 | 0.999 | 0.997 |
+| Spearman | 1.000 | 1.000 | 1.000 |
+| Mean | 2.03 | 1.64 | 2.10 |
+| Ranking Agreement | 4/4 | 4/4 | 4/4 |
+Appendix C Cross-Judge Evaluation
+To address the potential circularity of using Qwen3-Max as both the training reward provider and evaluation judge, we re-evaluate all model configurations using Gemini-2.5-Flash Google DeepMind (2025) as an independent judge with identical evaluation prompts.
+Gemini-2.5-Flash is not involved in any stage of CRAFT training.
+Table 12 reports the faithfulness scores from both judges across all model configurations under CRAFT.
+Gemini-2.5-Flash assigns systematically lower absolute scores than Qwen3-Max (mean = 1.92), which is expected given known differences in evaluation stringency across models.
+Crucially, the relative ordering of all four configurations is fully consistent under both judges across all three benchmarks (Spearman = 1.0).
+Table 13 confirms strong agreement: Pearson and Spearman across all benchmarks, demonstrating that the reported faithfulness improvements are robust to judge selection.
+Appendix D Case Studies
+| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
+[…]
+Subq1: Where is the Oh Yeah album from?
+Subq2: What is the largest city and capital of that state? Subq3: Who won the Indy car race in that city? |
+Subq1: Where is the Oh Yeah performer from?
+Subq2: What is the largest city of that state? Subq3: Who won the Indy car race there? |
+[…]
+[…]
+[…]
+[…]
+| 1 - Plan steps followed in order | 1 - Plan followed | |
+| 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs in gold_docs | |
+| 1 - Answer matches reasoning conclusion | 1 - Mario Andretti supported | |
+| 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified | |
+| 0.50 | 1.0 | |
+We present an additional case from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs.
+This case complements the right-answer-wrong-reason example in §4.6 by showing a case where both accuracy and faithfulness are corrected.
+D.1 Case: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
+Question: Who won the Indy car race in the largest city of the state where the Oh Yeah album performer was born?
+Gold Answer: Mario Andretti